### v1

# Equal contribution. * Corresponding author

Abstract
In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting and obstructions (e.g., handwear). The benchmark includes a diverse dataset from 28 subjects performing hand-object and hand-virtual interactions, accurately annotated with 3D hand poses through an automated process. We introduce a bespoken baseline method, TheFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery. Our experimental results highlight TheFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
1 Introduction
Egocentric 3D hand pose estimation is critically important for interpreting hand gestures across various applications, ranging from extended reality (XR) [93, 69, 57, 48], to human-robot interaction [71, 21, 22], and to imitation learning [66, 13, 1]. Its importance has been magnified with the advent of advanced XR headsets such as the Meta Quest series [58] and Apple Vision Pro [2], where it serves as a cornerstone for spatial interaction and immersive digital experiences.
While current research of hand pose estimation primarily focuses on RGB image-based methods [49, 47, 92, 20, 15], these approaches are particularly vulnerable to issues related to lighting variation and occlusions caused by handwear, e.g., gloves or large jewellery [86, 3]. These challenges underscore the imperative for robust egocentric 3D hand pose estimation capable of performing reliably in a variety of common yet complex daily scenarios. The prevailing approach to facilitate robust hand pose estimation in low-light conditions utilizes near infrared (NIR) cameras paired with active NIR emitters. This technology, invisible to the human eye, leverages active NIR emitter-receiver configurations for depth estimation through time-of-flight (ToF) or structured lighting. Nevertheless, active NIR systems are more power-intensive compared to passive sensing technologies [33, 32] and are prone to interference from external NIR sources, such as sunlight [84] and other NIR-equipped devices [75]. Consequently, these vulnerabilities restrict the effectiveness of hand pose estimation under bright daylight conditions and in situations where multiple augmented reality (AR) or virtual reality (VR) systems are used for collaborative works.
In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [54]. This unique attribute of thermal imaging introduces several benefits for 3D hand pose estimation. Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability. Moreover, thermal cameras are capable of detecting hands even under handwear such as gloves by identifying heat transmission patterns. This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cf. Tab. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions (cf. Fig. 1). To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cf. Tab. 2). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [68] (cf. Fig. 3).
Together with the multi-spectral dataset, we introduce a new baseline method named TheFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cf. Fig. 4). This approach is notable for its two parallel transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency. Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cf. Tab. 3). We then benchmark TheFormer against leading methods (cf. Tab. 4) and compare the performance of various spectral images (cf. Tab. 5 and Fig. 6). The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques. Our main contributions are summarized as follows:
  • $\bullet$
    We introduce the first-of-its-kind benchmark, dubbed ThermoHands, to investigate the potential of thermal imaging for egocentric 3D hand pose estimation.
  • $\bullet$
    We collected a diverse dataset comprising approximately 96,000 synchronized multi-spectral, multi-view images capturing hand-object and hand-virtual interactions from 28 participants across various environments. This dataset is enriched with 3D hand pose ground truths through an innovative automatic annotation process.
  • $\bullet$
    We introduce a new baseline method, termed TheFormer, and implement two state-of-the-art image-based methods on our dataset for benchmarking.
  • $\bullet$
    Based on the ThermoHands benchmark, we conduct comprehensive experiments and analysis on TheFormer and state-of-the-art methods.
  • $\bullet$
    We will release our dataset, code and models and maintain the benchmark to serve as a new challenge in 3D hand pose estimation.
2 Related Works
2.1 3D Hand Pose Datasets
Datasets with 3D hand pose annotations are imperative for training and evaluating ad-hoc models. Existing datasets, according to their approaches of annotation acquisition, can be summarized as four types in general, i.e., marker-based [97, 23, 85, 18], synthetic [63, 29, 62, 61, 103], manual [63, 83] or hybrid [11, 104, 60, 52, 64], and automatic [27, 7, 45, 76] annotated datasets. Marker-based approaches, using magnetic sensor [97, 23] or Mocap markers [85, 18], can alter and induce bias to the hand appearance. Synthetic data [63, 29, 61, 103, 62] suffers from the sim2real gap in terms of hand motion and texture features. Introducing human annotators to fully [63, 83] or partly [11, 104, 60, 52, 64] annotate 2D/3D keypoints circumvents the issues above, but it either limits the scale of datasets or manifests costly and laborious in practice. Most similar to ours, some datasets adopt fully automatic pipelines to obtain 3D hand pose annotations [27, 7, 45], which leverage pre-trained models (e.g. OpenPose [10]) to infer the prior hand information and rely on optimization to fit the MANO hand model [68].
Despite the existing progress, previous datasets only provide depth [97, 62], RGB images [61, 60, 104, 64] or both of them [23, 85, 29, 103, 63, 83, 11, 52, 27, 7, 45] as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation. ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images. Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications. Tab. 1 shows a comprehensive comparison between the existing datasets (in chronological order) and ours.
Table 1: Comparison of ThermoHands with existing image-based hand datasets with 3D pose annotations. (*) The number of synchronized frames of images. (**) No markers or sensors are attached to hands during data creation or capture.
Dataset (*)frames (**)markless depth thermal real ego exo two-hand hand-obj int hand-virt int
BigHands2.2M [97] 2.2M
SynthHands [63] 220K
EgoDexter [63] 3K
FPHA [23] 105K
ObMan [29] 150K
FreiHAND [104] 37K
HO-3D [27] 78K
ContactPose [7] 2.9M
InterHand2.6M [60] 2.6M
H2O [45] 571K
DexYCB [11] 508K
HOI4D [52] 2.4M
AssemblyHands [64] 3.0M
ARCTIC [18] 2.1M
ThermoHands (ours) 96K
2.2 Image-based 3D Hand Pose Estimation
As a key computer vision task, 3D hand pose estimation from images is highly demanded by applications like XR [93, 69, 57, 48], human-robot interaction [71, 21, 22] and imitation learning [66, 13, 1]. Therefore, this field has been extensively explored in previous arts that uses single RGB [35, 82, 96, 14, 39, 49, 60, 47, 80, 91, 98, 29, 51, 6, 4, 99] or depth [24, 59, 94, 100, 101] image as input. These methods can be roughly categorized into two fashions, i.e., model-based and model-free methods. Model-based methods [47, 80, 91, 98, 29, 51, 6, 4, 99] utilize the prior knowledge of the MANO hand model [68] by estimating its shape and pose parameters, while model-free methods [35, 82, 96, 14, 39, 49, 60, 24, 59, 94, 100, 101] learn the direct regression of 3D hand joints or vertices coordinates. Recently there has been growing interest in leveraging the temporal supervision [28, 51, 95, 43, 65] or leveraging sequential images as input [92, 20, 15, 9, 37, 16] for 3D hand pose estimation. In this study, we evaluate existing methods and our baseline method in both single image-based and video-based problem settings, respectively. Apart from the previous approaches, we investigate the potential of thermal imagery for tackling various challenges in 3D hand pose estimation.
2.3 Thermal Computer Vision
Thermal cameras achieve imaging by capturing the radiation emitted in the LWIR spectrum and deducing the temperature distribution on the surfaces [54]. Leveraging its robustness to variable illumination and unique temperature information, numerous efforts have been made to address various computer vision tasks, including super-resolution [67, 25, 36], human detection [8, 34, 26], action recognition [5, 17] and pose estimation [56, 12, 81], semantic segmentation [44, 41, 89, 46], depth estimation [78, 55, 77, 40], visual(-inertial) odometry/SLAM [38, 73, 79, 72], 3D reconstruction [50, 70, 74], etc. In this work, we focus on 3D hand pose estimation, which is an under-exploited task based on thermal images.
3 The ThermoHands Benchmark
3.1 Multi-Spectral Hand Pose Dataset
Figure 1: Data capture setup with the HMSP and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cf. Tab. 2), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the well-illuminated office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
Sensor Platforms. The HMSP is mounted with an Intel RealSense L515 LiDAR depth camera [33] streaming egocentric RGB, depth and NIR images, and a FLIR Boson 640 thermal camera [87] receiving the LWIR to obtain the thermal images. To better simulate XR devices, the HMSP is modularly designed as two components: a base and a sensor board with a fixed 30-degree downward tilt. An extra exocentric platform equipped with an Intel RealSense D455 [32] is leveraged to support multi-view annotation (cf. Sec. 3.2) as well as provide the RGB-D image data from the third-person viewpoint. As exhibited in Fig. 1, we place the two depth sensors outside each other’s FoV to minimize interference caused by their NIR emitters [33, 32]. In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
Synchronization. We use a single PC to simultaneously gather data streams from two sensor platforms, ensuring the synchronization of their timestamps. After collection, we synchronize six types of images, each with distinct frame rates, w.r.t. the timestamps of thermal images (8.5fps), thereby generating synchronized multi-spectral, multi-view data samples as our released data.
Table 2: Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
Well-illuminated office (Main) Other settings Total
Setting train val test sum darkness sun glare gloves kitchen
#frames 47,436 12,914 24,002 84,352 3,188 2,508 3,068 2,808 95,924
#seqs 172 43 86 301 12 12 12 14 352
#subjects 16 4 8 28 1 1 1 2 -
Figure 2: Thermal calibration chessboard containing a black base board and multiple removable white cubes (a). By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
Calibration. For accuracy, we use factory-calibrated intrinsic parameters from the D455 and L515 cameras [33, 32]. To better calibrate the parameters of the thermal camera, we self-design a modular calibration chessboard as shown in Fig. 2. Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig. 2. In this way, we can simultaneously calibrate the intrinsic parameter of the thermal camera and its extrinsic parameter w.r.t. the L515 camera. To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table. At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP [19]. For the subsequent frames, the KISS-ICP [90] odometry method is used to track the motion of the ego-head using the point clouds converted from the egocentric depth images.
Dataset Statistics. As shown in Tab. 2, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cf. Fig. 1) and 352 independent sequences in total. The main part is collected under the well-illuminated office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions. performs 7 scenario-specific hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands. This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2. We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen environment. Please refer to the supplementary for more dataset details.
Figure 3: Automatic annotation pipeline of 3D hand pose. We utilize the multi-view RGB and depth images as the input source and retrieve both 2D and 3D constraint information. Various error terms are formulated to optimize the MANO parameters.
3.2 Hand Pose Annotation
To refrain from employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [27, 7, 45], to obtain the 3D hand pose ground truth for our dataset.
In particular, we use the MANO statistical hand model [68] to represent 3D hand pose. The MANO model parameterizes the hand mesh vertices $\mathcal{V}(\beta,\theta)$ into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$, consisting of 45 parameters accounting for 15 hand joint angles (3 DoF for each) plus the rest for global rotation and translation. Following the MANO PyTorch version in [30], we denote the $N_{\mathcal{J}}=21$ hand joints mapped from the hand parameters as $\mathcal{J}(\beta,\theta)$. The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$ (1)
where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 3.
Initialization. For each sequence, we optimize the shape parameter $\beta$ [^4] For simplicity, we omit $\beta$ in Eq. 1 and subsequent equations. together with $\theta$ only at the first frame using Eq. 1 until convergence and fix its values for the subsequent frames. We initialize the pose parameter $\theta$ for each frame using the optimization result from the last frame. This helps to accelerate the convergence as well as keep the temporal consistency.
2D Joint Error $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}$. Given RGB images from $N_{C}$ viewpoints, we infer 2D hand joints $\mathcal{J}^{2D}$ using MediaPipe Hands [53] and define the 2D joint error as:
$\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{2D}_{c,i}-\pi_{c}(\mathcal{J}(\theta)_{i})||$ (2)
where $\pi_{c}(\cdot)$ returns the 2D projection location for 3D position in the $c$-th camera viewpoint, and $\alpha_{c}$ is hyperparameter used to weigh different viewpoints.
2D Mask Error $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}$. To generate the high-quality 2D hand mask, we prompt the prevalent Segment Anything Model [42] with the 2D hand joints $\mathcal{J}^{2D}$ and the bounding box derived from it. We penalize the distance between the hand mesh vertices $\mathcal{V}(\theta)$ and the 2D binary hand mask $\mathcal{M}_{c}$ as:
$\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{M}_{c,j}-\pi_{c}(\mathcal{V}(\theta)_{i})||$ (3)
where $\mathcal{M}_{c,j}$ is the coordinate of $j$-th non-zero pixel in the mask $\mathcal{M}_{c}$.
3D Joint Error $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$. We triangulate the 2D joints from multiple views to lift them to 3D joints $\mathcal{J}^{3D}$ and measure their difference to $\mathcal{J}(\theta)$, which is written as:
$\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}=\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{3D}_{i}-\mathcal{J}(\theta)_{i}||$ (4)
3D Mesh Error $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}$. To better supervise the hand mesh and fasten the optimization, we generate the 3D hand pose cloud $\mathcal{P}$ by cropping the depth image using the 2D hand mask and merging all views together. The 3D mesh error term compensates for the distance between the hand mesh and point cloud:
$\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}=\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{P}_{j}-\mathcal{V}(\theta)_{i}||$ (5)
Regularization $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$. To alleviate irregular hand articulation, we constrain the joint angles to pre-defined lower and upper boundaries $\underline{\theta}$ and $\overline{\theta}$:
$\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}=\sum_{i=1}^{45}(\max(\underline{\theta_{i}}-\theta_{i},0)+\max(\theta_{i}-\overline{\theta_{i}},0))$ (6)
Note that we also impose regularization to the shape parameter $\beta$ during initialization. Please see details of our annotation in supplementary material.
Figure 4: Overall Framework of TheFormer. Given thermal images, we extract both low-level and high-level features for each frame, which are input to the mask-guided spatial transformer and temporal transformer. Frame-wise spatial and temporal embeddings are concatenated and fed into the pose head to regress the 3D hand pose.
3.3 TheFormer: A Baseline Method
Our proposed baseline method, dubbed TheFormer, learns 3D hand pose estimation from egocentric thermal images based on Transformer, as exhibited in Fig. 4. The baseline network features in its two parallel transformer modules, i.e., mask-guided spatial transformer and temporal transformer, to model the spatio-temporal relationship of hand joints while being computationally efficient.
Problem Definition. We consider two problem settings: single image-based and video-based egocentric 3D hand pose estimation for our benchmark. In the former setting, we aim to estimate the 3D joint positions $\mathcal{J}_{t}$ for two hands given the single thermal image $\mathcal{I}_{t}$ captured for the $t$-th frame. For the video-based one, our input is a sequence of thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ and we estimate the per-frame 3D hand joint positions $\mathcal{J}_{i}$ together. Compared to the single image-based counterpart, the video-based setting allows for a fully temporal interaction among sequential images, potentially offering a better performance. Nevertheless, the accumulation of sequential images leads to a higher latency for online inference, limiting its usage to applications that demand real-time hand tracking.
Network Architecture. Without losing the generality, here we illustrate our network architecture for the video-based setting. Note that our network can be flexibly adapted to the single image-based problem by setting $T=1$. As seen in Fig. 4, given sequential thermal images, our architecture extracts the multi-level 2D feature using the backbone network per frame. We design a mask-guided spatial transformer module to enhance the spatial representation while enforcing the network to focus on the hand area. In parallel, we apply a temporal transformer module to reason the feature interaction along the temporal dimension. Both two transformers output per-frame feature embeddings that encode the spatial and temporal information, respectively. We concatenate them and use a pose head to regress the 3D coordinates of hand joints for two hands.
Backbone Features. For efficiency, we use the lightweight ResNet-18 [31] as our backbone network to extract multi-level features from thermal images. Specifically, we reserve both the low-level and high-level image features and input them to the spatial and temporal transformer modules individually. Low-level features carry fine-grained spatial details, which is crucial for accurate mask generation and spatial reasoning in our mask-guided spatial transformer. High-level features provide a broader context, helping temporal transformers to better understand the overall scene dynamics and the hand’s role within it.
Mask-guided Spatial Transformer. Human hands are highly articulated objects that can adopt a wide variety of poses, often against complex backgrounds. We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction. Given low-level features, we first utilize a mask head to estimate the binary hand mask in the thermal image. Then, we leverage the deformable self-attention [102] to refine the hand spatial features under the guidance of the hand mask. Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations. In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter. Lastly, we reduce the spatial dimensions of the spatial features via several downsample convolutional layers and flatten it into 1D spatial feature embedding per frame.
Temporal Transformer. Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities. To efficiently model temporal relationships, we apply the average-pooling to the high-level features to get the global feature vector per frame. The self-attention [88] is then employed to explicitly attend to the feature vector of every frame for a comprehensive temporal interaction. Similar to the spatial transformer, the output is frame-wise feature embedding that encompasses the temporal information.
Pose Regression. We concatenate the frame-wise spatial and temporal embedding from two transformer modules and get the spatio-temporal representations for each frame. In the pose head, we use the MLP to project the representations to the output space and obtain the per-frame 3D joint output $\mathcal{J}_{i}$.
Training Loss. We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cf. Sec. 3.2). For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately.
For the video-based setting, our baseline method is named as TheFormer-V, which accepts sequential images as input. The variant of our baseline to the single image-based setting is called TheFormer-S, where the temporal transformer degrades to a global transformer that only encodes the global information for a single frame. Please see the supplementary materials for more parameters, design and training details of our baseline method.
4 Experiment
4.1 Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset: one involving hand-object interaction and the other hand-virtual interaction, with a total of over 600 frames. To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
As shown in Tab. 3, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [45, 27, 104]. The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture. We also observe that only combining $\mathcal{L}_{mask}$ and $\mathcal{L}_{j2d}$ can already provide a plausible accuracy since they fit the projection of the 3D hand pose to two heterogeneous views. $\mathcal{L}_{mesh}$, though it hardly improves the joint accuracy, can result in more natural hand mesh. Adding $\mathcal{L}_{j3d}$ further refines the joints as it induces the explicit constraint to their positions. We showcase some annotation examples in Fig. 5. As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subject ethnicities.
Table 3: Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
Ego-view optimization Multi-view optimization
Errors $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}+\mathcal{L}_{j3d}$
mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation (defined by MANO [68]).
4.2 Benchmark Setup
Dataset Preparation. We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments. We annotate the main part of our dataset (cf. Tab. 1) in an automatic way following Sec. 3.2, of which the training and validation sets serve as the foundation for the training of all network models.
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation, i.e., HTT [92] and A2J-Transformer [35], and reproduce them on our dataset for benchmark experiments. HTT [92] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [35] only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TheFormer-V baseline. We also exclude the additional action block used by HTT [92] for hand action prediction, focusing solely on hand pose estimation. We adjusted the anchor initialization phase of the A2J-Transformer [35] to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU for a fair comparison of their inference speed.
Evaluation Metrics. We evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) [103], in both camera space and root-aligned (RA) space [60]. For RA space, we align the estimated wrist with its groundtruth position before measurement. For PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
4.3 Benchmark Results
Table 4: Benchmark results of TheFormer and state-of-the-art methods on thermal-based 3D hand pose estimation. Results are reported on the main testing set. $\uparrow$ denotes larger values are better, and vice versa. The fps displayed for sequence input indicates the number of sequences that models can process per second.
Method Input MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$ fps $\uparrow$
A2J-Transformer [35] Single 51.68 0.474 20.76 0.603 34
HTT [92] Single 49.09 0.489 20.69 0.599 211
TheFormer-S Single 48.25 0.510 22.60 0.565 120
HTT [92] Sequence 47.07 0.512 17.49 0.659 129
TheFormer-V Sequence 46.57 0.519 19.62 0.619 67
Comparison against State of the Arts. We compare the TheFormer against the state-of-the-art methods for thermal-based 3D hand pose estimation, as shown in Tab. 4. On the MEPE and AUC metrics, TheFormer-S outforms two competing methods under the single image-based setting, while TheFormer-V surpasses the counterpart HTT [92] given the same sequential images as input. Such an improvement stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cf. Sec. 3.3). A performance gap can be seen between single image-based and video-based settings for both HTT [92] and TheFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. We can also observe that TheFormer has slightly inferior results in the RA space relative to its opponents, leaving room for improvement in localizing individual hand joints once hand roots are correctly identified. This also suggests that our advancements are primarily attributed to our accurate localization of the overall hand especially the wrists. Thanks to our lightweight network design, TheFormer is highly efficient to run with fps of 120 and 67 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices.
Table 5: Comparison between different spectra. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
Method Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$
HTT - Sequence [92] RGB 43.30 0.542 15.43 0.697
Depth 41.62 0.559 17.66 0.654
NIR 41.57 0.562 16.41 0.679
Thermal 46.57 0.519 17.49 0.659
A2J-Transformer - Single [35] RGB 40.21 0.577 16.77 0.675
Depth 43.80 0.552 18.26 0.647
NIR 46.46 0.513 18.00 0.648
Thermal 51.68 0.474 20.76 0.603
Comparison between Spectrum. To compare the efficacy of different spectra (i.e., RGB, depth, NIR and thermal), we assess their performance by applying them as varying inputs to the HTT [92] and A2J-Transformer [35] method. As seen in Tab. 5, RGB spectrum yields favourable performance for both two methods as they provide rich multi-channel semantics, allowing for accurate hand pose reasoning. Specifically, A2J-Transformer [35] achieves comparable performance to HTT [92] with sequential input. This can be attributed to its usage of the larger ResNet50 [31] backbone that is pretrained with RGB images. However, this also increases the computation overhead, making A2J-Transformer [35] 6$\times$ slower than HTT [92] (cf. Tab. 4). Depth and NIR spectra show a marginal decrease in performance compared to RGB in well-illuminated conditions. However, the active illumination design of NIR sensors enables them to serve as supplements to RGB in dark conditions, as illustrated in Fig. 6. The thermal spectrum, operating independently of any active light source, still achieves plausible hand pose estimation accuracy, with a small discrepancy compared to other spectra. In addition to the above evaluation on the main testing set, we show more comparison of different spectra to highlight the value of thermal images under challenging conditions for 3D hand pose estimation.
4.4 Performance under Challenging Conditions
Figure 6: Examples of hand pose estimation results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on corresponding images.
To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison of four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf. Tab. 2). Our automatic annotation pipeline (cf. Sec. 3.2) becomes infeasible since hands appear corrupted in either RGB or depth images. Therefore, we conduct a qualitative analysis of their performance and present some representative examples in Fig. 6. As can be seen, RGB-based methods fail with gloves wearing (cf. Fig. 6 (a-b)) and in darkness (cf. Fig. 6 (c-d)). Gloves change how human hands look and hide their natural colors and textures. Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process. Contrary to RGB techniques, as shown in Fig. 6 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
NIR sensors are significantly disrupted by strong sunlight, affecting both NIR imaging and depth map creation, as shown in Fig. 6 (e-f). Conversely, thermal imaging is immune to sunlight and outdoor conditions. The temperature difference between hands and their surroundings in the thermal spectrum facilitates effortless identification of the hands, leaving the thermal-based estimator unaffected. The thermal camera’s ability to consistently capture hand features in diverse lighting conditions positions it as a suitable option for future XR applications. Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
5 Limitation and Future Work
Despite the advancements, we acknowledge certain limitations here guiding our future work. First, the scale of our current dataset is small when compared to some large-scale datasets [18, 64, 52, 60] that contain millions of images (cf. Tab. 1). We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions to better support future investigations in this field. Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other computer vision tasks relevant to human hands. In this regard, further efforts could be directed towards a) annotating the fine-grained hand action splits inside each sequence, b) leveraging the heat residual on objects after hand grasp to annotate the contact heatmap [7] and c) generating the ground truth mesh for both rigid [45] and articulated objects [18].
6 Conclusion
This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images. It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships. We demonstrate near 1cm annotation accuracy, show that TheFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios. We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.

### v3

@@
-Equal contribution. * Corresponding author
+ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images
 Abstract
-In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting and obstructions (e.g., handwear).
-The benchmark includes a diverse dataset from 28 subjects performing hand-object and hand-virtual interactions, accurately annotated with 3D hand poses through an automated process.
-We introduce a bespoken baseline method, TheFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery.
-Our experimental results highlight TheFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
+In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting conditions and obstructions (e.g., handwear).
+The benchmark includes a multi-view and multi-spectral dataset collected from 28 subjects performing hand-object and hand-virtual interactions under diverse scenarios, accurately annotated with 3D hand poses through an automated process.
+We introduce a new baseline method, TherFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery.
+Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
 1 Introduction
@@
 In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [24].
-This unique attribute of thermal imaging introduces several benefits for 3D hand pose estimation.
+This unique attribute of thermal imaging introduces several benefits for hand pose estimation.
 Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability.
-Moreover, thermal cameras are capable of detecting hands even under handwear such as gloves by identifying heat transmission patterns.
+Moreover, thermal cameras can detect hands even under handwear such as gloves by identifying heat transmission patterns.
 This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
@@
 To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging.
-This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cf.
-Tab.
+This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig.
 1).
-Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions (cf.
-Fig.
+Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions.
+To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab.
 1).
-To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cf.
-Tab.
+Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline.
+This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [25] (cfFig.
+3).
+Together with the multi-spectral dataset, we introduce a new baseline method named TherFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cfFig.
+4).
+This approach is notable for its two consecutive transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency.
+Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cfTab.
 2).
-Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline.
-This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [68] (cf.
-Fig.
-3).
-Together with the multi-spectral dataset, we introduce a new baseline method named TheFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cf.
-Fig.
-4).
-This approach is notable for its two parallel transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency.
-Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cf.
-Tab.
-3).
-We then benchmark TheFormer against leading methods (cf.
-Tab.
-4) and compare the performance of various spectral images (cf.
-Tab.
+We then benchmark TherFormer against leading methods (cfTab.
+3) and compare the performance of various spectral images (cfTab.
 4 and Fig.
@@
   • $\bullet$
-    We introduce a new baseline method, termed TheFormer, and implement two state-of-the-art image-based methods on our dataset for benchmarking.
+    We introduce a new baseline method, termed TherFormer, and implement state-of-the-art image-based methods on our dataset for benchmarking.
   • $\bullet$
-    Based on the ThermoHands benchmark, we conduct comprehensive experiments and analysis on TheFormer and state-of-the-art methods.
+    Based on the ThermoHands benchmark, we conduct comprehensive experiments and analysis on TherFormer and state-of-the-art methods.
   • $\bullet$
@@
 Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications.
-Tab.
-1 shows a comprehensive comparison between the existing datasets (in chronological order) and ours.
-Table 1: Comparison of ThermoHands with existing image-based hand datasets with 3D pose annotations.
-(*) The number of synchronized frames of images.
-(**) No markers or sensors are attached to hands during data creation or capture.
-Dataset (*)frames (**)markless depth thermal real ego exo two-hand hand-obj int hand-virt int
-BigHands2.2M [97] 2.2M
-SynthHands [63] 220K
-EgoDexter [63] 3K
-FPHA [23] 105K
-ObMan [29] 150K
-FreiHAND [104] 37K
-HO-3D [27] 78K
-ContactPose [7] 2.9M
-InterHand2.6M [60] 2.6M
-H2O [45] 571K
-DexYCB [11] 508K
-HOI4D [52] 2.4M
-AssemblyHands [64] 3.0M
-ARCTIC [18] 2.1M
-ThermoHands (ours) 96K
+Figure 1: Data capture setup with the HMSP and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
 2.2 Image-based 3D Hand Pose Estimation
@@
 3.1 Multi-Spectral Hand Pose Dataset
-Figure 1: Data capture setup with the HMSP and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
 Overview.
-At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cf.
-Tab.
-2), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection..
+At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab.
+1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from School of Informatics, University of Edinburgh, and participant consent forms were signed before the collection..
 As shown in Fig.
 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data.
-During capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table.
-The main part is captured in the well-illuminated office scenario.
+During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table.
+The main part is captured in the normal office scenario.
 To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
+Normal office (Main) Other settings Total
+Setting train val test sum darkness sun glare gloves kitchen
+#frames 47,436 12,914 24,002 84,352 3,188 2,508 3,068 2,808 95,924
+#seqs 172 43 86 301 12 12 12 14 352
+#subjects 16 4 8 28 1 1 1 2 -
+Table 1: Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
+Figure 2: Thermal calibration chessboard containing a black base board and multiple removable white cubes (a).
+By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
 Sensor Platforms.
@@
 As exhibited in Fig.
-1, we place the two depth sensors outside each other’s FoV to minimize interference caused by their NIR emitters [33, 32].
+1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters [20, 21].
 In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
@@
 After collection, we synchronize six types of images, each with distinct frame rates, w.r.t.
-the timestamps of thermal images (8.5fps), thereby generating synchronized multi-spectral, multi-view data samples as our released data.
-Table 2: Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
-Well-illuminated office (Main) Other settings Total
-Setting train val test sum darkness sun glare gloves kitchen
-#frames 47,436 12,914 24,002 84,352 3,188 2,508 3,068 2,808 95,924
-#seqs 172 43 86 301 12 12 12 14 352
-#subjects 16 4 8 28 1 1 1 2 -
-Figure 2: Thermal calibration chessboard containing a black base board and multiple removable white cubes (a).
-By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
+the timestamps of thermal images (8.5fps), thereby generating well-aligned multi-spectral, multi-view data samples as our released data.
 Calibration.
@@
 2.
-In this way, we can simultaneously calibrate the intrinsic parameter of the thermal camera and its extrinsic parameter w.r.t.
-the L515 camera.
+In this way, we can simultaneously calibrate the intrinsic parameter of the L515 thermal camera and its extrinsic parameter w.r.t.
+the D455 RGB-D camera.
 To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table.
-At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP [19].
-For the subsequent frames, the KISS-ICP [90] odometry method is used to track the motion of the ego-head using the point clouds converted from the egocentric depth images.
+At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP problem [97].
+For the subsequent frames, as the exocentric platform keeps stationary, we only track the the motion of the ego-head using KISS-ICP [98] odometry method, which takes the point clouds converted from the egocentric depth image as input.
 Dataset Statistics.
-As shown in Tab.
-2, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cf.
-Fig.
+As seen in Tab.
+1, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cfFig.
 1) and 352 independent sequences in total.
-The main part is collected under the well-illuminated office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions.
-performs 7 scenario-specific hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands.
+The main part is collected under the normal office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions.
+performs 7 hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands.
 This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2.
-We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen environment.
+We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen.
 Please refer to the supplementary for more dataset details.
-Figure 3: Automatic annotation pipeline of 3D hand pose.
-We utilize the multi-view RGB and depth images as the input source and retrieve both 2D and 3D constraint information.
-Various error terms are formulated to optimize the MANO parameters.
 3.2 Hand Pose Annotation
-To refrain from employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [27, 7, 45], to obtain the 3D hand pose ground truth for our dataset.
+To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [41, 42, 43], to obtain the 3D hand pose ground truth for our dataset.
 In particular, we use the MANO statistical hand model [25] to represent 3D hand pose.
-The MANO model parameterizes the hand mesh vertices $\mathcal{V}(\beta,\theta)$ into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$, consisting of 45 parameters accounting for 15 hand joint angles (3 DoF for each) plus the rest for global rotation and translation.
-Following the MANO PyTorch version in [30], we denote the $N_{\mathcal{J}}=21$ hand joints mapped from the hand parameters as $\mathcal{J}(\beta,\theta)$.
+The MANO model parameterizes the hand mesh vertices into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$.
 The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
 $\theta^{*}=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$ (1)
-where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 3.
-Initialization.
-For each sequence, we optimize the shape parameter $\beta$ [^4] For simplicity, we omit $\beta$ in Eq.
-1 and subsequent equations.
-together with $\theta$ only at the first frame using Eq.
-1 until convergence and fix its values for the subsequent frames.
-We initialize the pose parameter $\theta$ for each frame using the optimization result from the last frame.
-This helps to accelerate the convergence as well as keep the temporal consistency.
-2D Joint Error $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}$.
-Given RGB images from $N_{C}$ viewpoints, we infer 2D hand joints $\mathcal{J}^{2D}$ using MediaPipe Hands [53] and define the 2D joint error as:
-$\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{2D}_{c,i}-\pi_{c}(\mathcal{J}(\theta)_{i})||$ (2)
-where $\pi_{c}(\cdot)$ returns the 2D projection location for 3D position in the $c$-th camera viewpoint, and $\alpha_{c}$ is hyperparameter used to weigh different viewpoints.
-2D Mask Error $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}$.
-To generate the high-quality 2D hand mask, we prompt the prevalent Segment Anything Model [42] with the 2D hand joints $\mathcal{J}^{2D}$ and the bounding box derived from it.
-We penalize the distance between the hand mesh vertices $\mathcal{V}(\theta)$ and the 2D binary hand mask $\mathcal{M}_{c}$ as:
-$\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{M}_{c,j}-\pi_{c}(\mathcal{V}(\theta)_{i})||$ (3)
-where $\mathcal{M}_{c,j}$ is the coordinate of $j$-th non-zero pixel in the mask $\mathcal{M}_{c}$.
-3D Joint Error $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$.
-We triangulate the 2D joints from multiple views to lift them to 3D joints $\mathcal{J}^{3D}$ and measure their difference to $\mathcal{J}(\theta)$, which is written as:
-$\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}=\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{3D}_{i}-\mathcal{J}(\theta)_{i}||$ (4)
-3D Mesh Error $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}$.
-To better supervise the hand mesh and fasten the optimization, we generate the 3D hand pose cloud $\mathcal{P}$ by cropping the depth image using the 2D hand mask and merging all views together.
-The 3D mesh error term compensates for the distance between the hand mesh and point cloud:
-$\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}=\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{P}_{j}-\mathcal{V}(\theta)_{i}||$ (5)
-Regularization $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$.
-To alleviate irregular hand articulation, we constrain the joint angles to pre-defined lower and upper boundaries $\underline{\theta}$ and $\overline{\theta}$:
-$\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}=\sum_{i=1}^{45}(\max(\underline{\theta_{i}}-\theta_{i},0)+\max(\theta_{i}-\overline{\theta_{i}},0))$ (6)
-Note that we also impose regularization to the shape parameter $\beta$ during initialization. Please see details of our annotation in supplementary material.
-Figure 4: Overall Framework of TheFormer.
-Given thermal images, we extract both low-level and high-level features for each frame, which are input to the mask-guided spatial transformer and temporal transformer.
-Frame-wise spatial and temporal embeddings are concatenated and fed into the pose head to regress the 3D hand pose.
-3.3 TheFormer: A Baseline Method
-Our proposed baseline method, dubbed TheFormer, learns 3D hand pose estimation from egocentric thermal images based on Transformer, as exhibited in Fig.
-4.
-The baseline network features in its two parallel transformer modules, i.e., mask-guided spatial transformer and temporal transformer, to model the spatio-temporal relationship of hand joints while being computationally efficient.
+where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors.
+The diagram illustrating our annotation process is shown in Fig.
+3.
+For more explanation of error terms and implementation details of our dataset annotation process, please refer to our supplementary document.
+Figure 3: Automatic annotation pipeline of 3D hand pose.
+We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands [99] and SAM [100].
+Various error terms are formulated to optimize the MANO parameters.
+3.3 TherFormer: A Baseline Method
+For our benchmark, we setup a baseline method, dubbed TherFormer, for thermal image-based 3D hand pose estimation.
+As exhibited in Fig.
+4, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient.
+Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images.
+However, our primary objective is to establish a baseline method for future research, rather than focusing exclusively on methods for thermal images.
+Figure 4: Overall Framework of TherFormer.
+Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction.
+Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
 Problem Definition.
@@
 For the video-based one, our input is a sequence of thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ and we estimate the per-frame 3D hand joint positions $\mathcal{J}_{i}$ together.
-Compared to the single image-based counterpart, the video-based setting allows for a fully temporal interaction among sequential images, potentially offering a better performance.
-Nevertheless, the accumulation of sequential images leads to a higher latency for online inference, limiting its usage to applications that demand real-time hand tracking.
-Network Architecture.
-Without losing the generality, here we illustrate our network architecture for the video-based setting.
-Note that our network can be flexibly adapted to the single image-based problem by setting $T=1$.
-As seen in Fig.
-4, given sequential thermal images, our architecture extracts the multi-level 2D feature using the backbone network per frame.
-We design a mask-guided spatial transformer module to enhance the spatial representation while enforcing the network to focus on the hand area.
-In parallel, we apply a temporal transformer module to reason the feature interaction along the temporal dimension.
-Both two transformers output per-frame feature embeddings that encode the spatial and temporal information, respectively.
-We concatenate them and use a pose head to regress the 3D coordinates of hand joints for two hands.
-Backbone Features.
-For efficiency, we use the lightweight ResNet-18 [31] as our backbone network to extract multi-level features from thermal images.
-Specifically, we reserve both the low-level and high-level image features and input them to the spatial and temporal transformer modules individually.
-Low-level features carry fine-grained spatial details, which is crucial for accurate mask generation and spatial reasoning in our mask-guided spatial transformer.
-High-level features provide a broader context, helping temporal transformers to better understand the overall scene dynamics and the hand’s role within it.
+Without losing the generality, here we illustrate our network architecture for the video-based setting, i.e., TherFormer-V.
+In practice, our network can be flexibly adapted to the single image-based version (i.e., TherFormer-S) by setting $T=1$.
 Mask-guided Spatial Transformer.
@@
 We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction.
-Given low-level features, we first utilize a mask head to estimate the binary hand mask in the thermal image.
-Then, we leverage the deformable self-attention [102] to refine the hand spatial features under the guidance of the hand mask.
+Given backbone features, we first utilize a mask head to estimate the binary hand mask in the thermal image.
+Then, we leverage the deformable self-attention [101] to refine the hand spatial features under the guidance of the estimated hand mask.
 Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations.
 In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter.
-Lastly, we reduce the spatial dimensions of the spatial features via several downsample convolutional layers and flatten it into 1D spatial feature embedding per frame.
+Lastly, we reduce the spatial dimensions of the spatial features with a series of convolutions layers for efficiency.
 Temporal Transformer.
 Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities.
-To efficiently model temporal relationships, we apply the average-pooling to the high-level features to get the global feature vector per frame.
-The self-attention [88] is then employed to explicitly attend to the feature vector of every frame for a comprehensive temporal interaction.
-Similar to the spatial transformer, the output is frame-wise feature embedding that encompasses the temporal information.
-Pose Regression.
-We concatenate the frame-wise spatial and temporal embedding from two transformer modules and get the spatio-temporal representations for each frame.
-In the pose head, we use the MLP to project the representations to the output space and obtain the per-frame 3D joint output $\mathcal{J}_{i}$.
-Training Loss.
-We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cf.
-Sec.
+To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention [102] to explicitly attend to the feature vector of every frame.
+The output is frame-wise spatio-temporal feature embeddings.
+Note that the temporal self-attention degrades to an MLP for single-image based setting.
+In the pose head, we use the MLP to project the embeddings to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$.
+We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cfSec.
 3.2).
 For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately.
-For the video-based setting, our baseline method is named as TheFormer-V, which accepts sequential images as input.
-The variant of our baseline to the single image-based setting is called TheFormer-S, where the temporal transformer degrades to a global transformer that only encodes the global information for a single frame.
-Please see the supplementary materials for more parameters, design and training details of our baseline method.
+Please see more details of TherFormer in our supplementary document.
 4 Experiment
 4.1 Evaluation of the Annotation Method
-As the first step, we validate the accuracy of our 3D hand pose annotation and analyze the impact for optimization results.
-For evaluation, we manually annotate two random sequences from our main dataset: one involving hand-object interaction and the other hand-virtual interaction, with a total of over 600 frames.
-To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D positions by triangulation.
+As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec.
+3.2) and analyze the impact for optimization results.
+For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames.
+To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D joint positions by triangulation.
 We calculate the average 3D joint errors across all frames to measure the accuracy.
 As shown in Tab.
-3, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [45, 27, 104].
+2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [43, 41, 37].
 The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture.
@@
 5.
-As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subject ethnicities.
-Table 3: Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
+As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subjects’ hand color and shape.
+Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
 Ego-view optimization Multi-view optimization
@@
 mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
-Figure 5: Examples of 3D hand pose annotations.
-Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images.
-Bottom row: visualization of hand mesh annotation (defined by MANO [68]).
-4.2 Benchmark Setup
+Table 2: Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
+4.2 Experiment Setup
 Dataset Preparation.
 We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments.
-We annotate the main part of our dataset (cf.
-Tab.
-1) in an automatic way following Sec.
+We annotate the main part of our dataset (cfTab.
+1) automatically following Sec.
 3.2, of which the training and validation sets serve as the foundation for the training of all network models.
 Method and Implementation.
-To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation, i.e., HTT [92] and A2J-Transformer [35], and reproduce them on our dataset for benchmark experiments.
+To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT [15] and A2J-Transformer [46], and reproduced them on our dataset for experiments.
 HTT [15] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [46] only works under the single image-based setting.
-For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TheFormer-V baseline.
-We also exclude the additional action block used by HTT [92] for hand action prediction, focusing solely on hand pose estimation.
-We adjusted the anchor initialization phase of the A2J-Transformer [35] to better accommodate our dataset, without altering the density of its anchors.
-All trained models are tested on a single NVIDIA RTX 4090 GPU for a fair comparison of their inference speed.
+For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines.
+We exclude the additional action block used by HTT [15], focusing solely on hand pose estimation.
+We adjusted the anchor initialization phase of A2J-Transformer [46] to better accommodate our dataset, without altering the density of its anchors.
+All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
 Evaluation Metrics.
@@
 For PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
-4.3 Benchmark Results
-Table 4: Benchmark results of TheFormer and state-of-the-art methods on thermal-based 3D hand pose estimation.
-Results are reported on the main testing set.
-$\uparrow$ denotes larger values are better, and vice versa.
-The fps displayed for sequence input indicates the number of sequences that models can process per second.
+4.3 Thermal Image-based 3D Hand Pose Estimation
+Main Results.
+To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods [15, 46] on the main testing set, as shown in Tab.
+3.
+TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT [15] given the same sequential images as input.
+Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec.
+3.3).
+A performance gap can be observed between single image-based and video-based settings for both HTT [15] and TherFormer.
+We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities.
+Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices.
+Moreover, TherFormer-V also improves the performance over HTT [15] for other spectra (cfTab.
+4), proving its adaptability to different inputs.
 Method Input MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$ fps $\uparrow$
-A2J-Transformer [35] Single 51.68 0.474 20.76 0.603 34
-HTT [92] Single 49.09 0.489 20.69 0.599 211
-TheFormer-S Single 48.25 0.510 22.60 0.565 120
-HTT [92] Sequence 47.07 0.512 17.49 0.659 129
-TheFormer-V Sequence 46.57 0.519 19.62 0.619 67
-Comparison against State of the Arts.
-We compare the TheFormer against the state-of-the-art methods for thermal-based 3D hand pose estimation, as shown in Tab.
-4.
-On the MEPE and AUC metrics, TheFormer-S outforms two competing methods under the single image-based setting, while TheFormer-V surpasses the counterpart HTT [92] given the same sequential images as input.
-Such an improvement stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cf.
-Sec.
-3.3).
-A performance gap can be seen between single image-based and video-based settings for both HTT [92] and TheFormer.
-We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities.
-We can also observe that TheFormer has slightly inferior results in the RA space relative to its opponents, leaving room for improvement in localizing individual hand joints once hand roots are correctly identified.
-This also suggests that our advancements are primarily attributed to our accurate localization of the overall hand especially the wrists.
-Thanks to our lightweight network design, TheFormer is highly efficient to run with fps of 120 and 67 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices.
-Table 5: Comparison between different spectra. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
-Method Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$
-HTT - Sequence [92] RGB 43.30 0.542 15.43 0.697
-Depth 41.62 0.559 17.66 0.654
-NIR 41.57 0.562 16.41 0.679
-Thermal 46.57 0.519 17.49 0.659
-A2J-Transformer - Single [35] RGB 40.21 0.577 16.77 0.675
-Depth 43.80 0.552 18.26 0.647
-NIR 46.46 0.513 18.00 0.648
-Thermal 51.68 0.474 20.76 0.603
-Comparison between Spectrum.
-To compare the efficacy of different spectra (i.e., RGB, depth, NIR and thermal), we assess their performance by applying them as varying inputs to the HTT [92] and A2J-Transformer [35] method.
-As seen in Tab.
-5, RGB spectrum yields favourable performance for both two methods as they provide rich multi-channel semantics, allowing for accurate hand pose reasoning.
-Specifically, A2J-Transformer [35] achieves comparable performance to HTT [92] with sequential input.
-This can be attributed to its usage of the larger ResNet50 [31] backbone that is pretrained with RGB images.
-However, this also increases the computation overhead, making A2J-Transformer [35] 6$\times$ slower than HTT [92] (cf.
+(a) A2J-Transformer [46] Single 51.68 0.474 20.76 0.603 34
+(b) HTT [15] Single 49.09 0.489 20.69 0.599 211
+(c) TherFormer-S Single 44.64 0.539 18.34 0.643 136
+(d) (c) w/o spatial transformer Single 48.79 0.491 20.15 0.609 174
+(e) (c) w/o mask guidance Single 48.83 0.494 18.89 0.625 141
+(f) HTT [15] Sequence 47.07 0.512 17.49 0.659 129
+(g) TherFormer-V Sequence 43.36 0.549 17.36 0.661 52
+Table 3: Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
+Ablation Studies.
+We ablate our proposed mask-guided attention mechanism and the entire spatial transformer (cfSec.
+3.3) to observe their impact.
+As can be seen in Tab.
+3, the mask-guided attention mechanism notably contributes to TherFormer’s performance growth (row (c) vs.
+(e)), demonstrating its effectiveness in mitigating the effect of background clutter.
+Other components in the spatial transformer bring a large performance gain in the RA space (row (e) vs.
+(d)), confirming the importance of spatial feature enhancement for fine-grained hand joints localization.
+4.4 Comparison between Spectrum
+4.4.1 Qualitative Results under Challenging Conditions
+To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf.
 Tab.
-4).
-Depth and NIR spectra show a marginal decrease in performance compared to RGB in well-illuminated conditions.
-However, the active illumination design of NIR sensors enables them to serve as supplements to RGB in dark conditions, as illustrated in Fig.
-6.
-The thermal spectrum, operating independently of any active light source, still achieves plausible hand pose estimation accuracy, with a small discrepancy compared to other spectra.
-In addition to the above evaluation on the main testing set, we show more comparison of different spectra to highlight the value of thermal images under challenging conditions for 3D hand pose estimation.
-4.4 Performance under Challenging Conditions
-Figure 6: Examples of hand pose estimation results with various spectra under three challenging settings.
-For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on corresponding images.
-To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison of four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf.
-Tab.
-2).
+1).
 Our automatic annotation pipeline (cf.
@@
 6.
-As can be seen, RGB-based methods fail with gloves wearing (cf.
-Fig.
-6 (a-b)) and in darkness (cf.
-Fig.
+We use the same model, i.e., HTT [15], trained on different spectra for a fair comparison.
+Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
+RGB vs.
+Thermal.
+As can be seen, RGB-based methods fail with gloves wearing (cfFig.
+6 (a-b)) and in darkness (cfFig.
 6 (c-d)).
@@
 6 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
+NIR, Depth vs.
+Thermal.
 NIR sensors are significantly disrupted by strong sunlight, affecting both NIR imaging and depth map creation, as shown in Fig.
@@
 The thermal camera’s ability to consistently capture hand features in diverse lighting conditions positions it as a suitable option for future XR applications.
-Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
+Figure 6: Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
+4.4.2 Quantitative Results under Normal Conditions
+To validate the versatility of thermal-based methods, we also evaluate their performance compared to other spectra (i.e., RGB, depth, NIR and thermal) under normal conditions for general-purpose use cases.
+We apply different spectral images as input to multiple baselines methods and report the results in Tab.
+4.
+Not surprisingly, depth image-based methods yield the best performance on average, as they directly utilize the depth information to provide a detailed 3D structure of the hand.
+The NIR spectrum shows a marginal decrease in performance compared to depth but outperforms both RGB and thermal spectra.
+This can be attributed to NIR camera’s active sensing ability, which leads to more consistent and reliable imaging regardless of the variability of external illuminations (vs.
+RGB) and temperature (vs.
+thermal).
+However, active NIR sensors are prone to interference from external NIR sources like sun glare (cfFig.
+6).
+Thermal-based methods, despite using single-channel heat capture, achieves comparable performance to those using RGB images that contain rich color and textures.
+Such results suggest that thermal images can not only serve as supplements in challenging cases but also can be a viable alternative to other spectra in normal conditions for hand pose estimation.
+HTT (Sequence) TherFormer-V TherFormer-S Average
+Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$
+RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.80 0.538
+Depth 41.62 0.559 39.70 0.581 39.84 0.579 40.39 0.573
+NIR 41.57 0.562 40.79 0.575 40.98 0.575 41.11 0.571
+Thermal 47.07 0.512 43.36 0.549 44.64 0.539 45.06 0.533
+Table 4: Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
 5 Limitation and Future Work
 Despite the advancements, we acknowledge certain limitations here guiding our future work.
-First, the scale of our current dataset is small when compared to some large-scale datasets [18, 64, 52, 60] that contain millions of images (cf.
-Tab.
-1).
-We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions to better support future investigations in this field.
-Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other computer vision tasks relevant to human hands.
-In this regard, further efforts could be directed towards a) annotating the fine-grained hand action splits inside each sequence, b) leveraging the heat residual on objects after hand grasp to annotate the contact heatmap [7] and c) generating the ground truth mesh for both rigid [45] and articulated objects [18].
+First, the scale of our current dataset is relatively small when compared to some datasets [29, 40, 39, 38].
+We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions.
+Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands.
+Further efforts will be annotating the fine-grained hand action splits and contact heatmap for hand grasp [42].
+Lastly, we currently rely on qualitative analysis for challenging conditions when our automatic annotation method becomes unreliable.
+In the future, we will leverage unaffected spectral images in individual cases to manually annotate the 3D hand pose.
 6 Conclusion
@@
 It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships.
-We demonstrate near 1cm annotation accuracy, show that TheFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios.
+We demonstrate near 1cm annotation accuracy, show that TherFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios.
 We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.
+Acknowledgments and Disclosure of Funding
+This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1), and partially supported by Cisco Research.
+Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+[Yes] We elaborate our contributions as a list in the introduction section and briefly explain them in the abstract.
+    2. (b)
+      Did you describe the limitations of your work? [Yes] See Sec. 5
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work? [N/A] There is no societal impact of this work.
+    4. (d)
+      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We have reviewed the ethics review guidelines and make sure we obey all rules for this research.
+  2. 2.
+    If you are including theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? [N/A] This work does not contribute to the theoretical results.
+    2. (b)
+      Did you include complete proofs of all theoretical results? [N/A] This work does not contribute to the theoretical results.
+  3. 3.
+    If you ran experiments (e.g. for benchmarks)…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
+[No] We will release our code upon the acceptance of this paper.
+    2. (b)
+      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We describe them in our main paper and supplementary document.
+    3. (c)
+      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
+[No] In line with the state-of-the-art 3D hand pose estimation works, we do not report error bars for our experiment results.
+    4. (d)
+      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We include the type of GPUs in the main paper.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+    1. (a)
+      If your work uses existing assets, did you cite the creators? [Yes] We use the offical code of HTT [15] and A2J-transformer [46] to reproduce their results on our dataset, which have been cited.
+    2. (b)
+      Did you mention the license of the assets? [Yes] We mention the license of the offical codes of HTT and A2J-Transformer in the supplementary document.
+    3. (c)
+      Did you include any new assets either in the supplemental material or as a URL? [Yes] We include the datasheet and the download URL of our dataset in the supplemental document.
+    4. (d)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
+[Yes] We claim in the footnote that participant consent forms were signed before the collection.
+    5. (e)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We discuss in the supplementary document.
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We do not give written instructions to our participants.
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] There is no potential participant risks.
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] Our participants are volunteer.

### v4

@@
 ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images
-Abstract
-In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting conditions and obstructions (e.g., handwear).
+Abstract.
+Designing egocentric 3D hand pose estimation systems that can perform reliably in complex, real-world scenarios is crucial for downstream applications.
+Previous approaches using RGB or NIR imagery struggle in challenging conditions: RGB methods are susceptible to lighting variations and obstructions like handwear, while NIR techniques can be disrupted by sunlight or interference from other NIR-equipped devices.
+To address these limitations, we present ThermoHands, the first benchmark focused on thermal image-based egocentric 3D hand pose estimation, demonstrating the potential of thermal imaging to achieve robust performance under these conditions.
 The benchmark includes a multi-view and multi-spectral dataset collected from 28 subjects performing hand-object and hand-virtual interactions under diverse scenarios, accurately annotated with 3D hand poses through an automated process.
@@
 Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
-1 Introduction
-Egocentric 3D hand pose estimation is critically important for interpreting hand gestures across various applications, ranging from extended reality (XR) [1, 2, 3, 4], to human-robot interaction [5, 6, 7], and to imitation learning [8, 9, 10].
-Its importance has been magnified with the advent of advanced XR headsets such as the Meta Quest series [11] and Apple Vision Pro [12], where it serves as a cornerstone for spatial interaction and immersive digital experiences.
-While current research of hand pose estimation primarily focuses on RGB image-based methods [13, 14, 15, 16, 17], these approaches are particularly vulnerable to issues related to lighting variation and occlusions caused by handwear, e.g., gloves or large jewellery [18, 19].
+* Equal contribution
+† Corresponding author: xiaoxuan.lu@ucl.ac.uk
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
+Figure 1. Data capture setup with the customized head-mounted sensor platform (HMSP) and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
+1. Introduction
+Egocentric 3D hand pose estimation is critically important for interpreting hand gestures across various applications, ranging from extended reality (XR) (wu2020hand, ; sagayam2017hand, ; marchand2015pose, ; liang2015ar, ), to human-robot interaction (sampieri2022pose, ; gao2021dynamic, ; gao2019dual, ), and to imitation learning (qin2022dexmv, ; chen2023bi, ; antotsiou2018task, ).
+Its importance has been magnified with the advent of advanced XR headsets such as the Meta Quest series (meta_quest, ) and Apple Vision Pro (apple_vision_pro_2024, ), where it serves as a cornerstone for spatial interaction and immersive digital experiences.
+While current research of hand pose estimation primarily focuses on RGB image-based methods (lin2021two, ; li2022interacting, ; wen2023hierarchical, ; Fu_2023_ICCV, ; cho2023transformer, ), these approaches are particularly vulnerable to issues related to lighting variation and occlusions caused by handwear, e.g., gloves or large jewellery (youtube_video_2024, ; apple_support_gloves_2024, ).
 These challenges underscore the imperative for robust egocentric 3D hand pose estimation capable of performing reliably in a variety of common yet complex daily scenarios.
@@
 This technology, invisible to the human eye, leverages active NIR emitter-receiver configurations for depth estimation through time-of-flight (ToF) or structured lighting.
-Nevertheless, active NIR systems are more power-intensive compared to passive sensing technologies [20, 21] and are prone to interference from external NIR sources, such as sunlight [22] and other NIR-equipped devices [23].
+Nevertheless, active NIR systems are more power-intensive compared to passive sensing technologies (IntelRSL515, ; IntelRealSense2023D455, ) and are prone to interference from external NIR sources, such as sunlight (suarez2012using, ) and other NIR-equipped devices (seewald2019toward, ).
 Consequently, these vulnerabilities restrict the effectiveness of hand pose estimation under bright daylight conditions and in situations where multiple augmented reality (AR) or virtual reality (VR) systems are used for collaborative works.
-In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [24].
+In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum (lloyd2013thermal, ).
 This unique attribute of thermal imaging introduces several benefits for hand pose estimation.
@@
 This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
-Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap.
+Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR images as two ‘spectra’, despite their usual overlap.
 spectral imagery?
-To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging.
+To answer this, we introduce ThermoHands  [^2] Anonymous project page: https://thermohands.github.io/, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging.
 This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig.
@@
 Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline.
-This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [25] (cfFig.
-3).
+This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model (MANO, ) (cfFig.
+4).
 Together with the multi-spectral dataset, we introduce a new baseline method named TherFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cfFig.
-4).
+5).
 This approach is notable for its two consecutive transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency.
@@
 3) and compare the performance of various spectral images (cfTab.
+5, Tab.
 4 and Fig.
-6).
+8).
 The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques.
@@
     We will release our dataset, code and models and maintain the benchmark to serve as a new challenge in 3D hand pose estimation.
-2 Related Works
-2.1 3D Hand Pose Datasets
+2. Related Works
+Normal office (Main) Other settings Total
+Setting train val test sum darkness sun glare gloves kitchen
+#frames 47,436 12,914 24,002 84,352 3,188 2,508 3,068 2,808 95,924
+#seqs 172 43 86 301 12 12 12 14 352
+#subjects 16 4 8 28 1 1 1 2 -
+Table 1. Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
+2.1. 3D Hand Pose Datasets
 Datasets with 3D hand pose annotations are imperative for training and evaluating ad-hoc models.
-Existing datasets, according to their approaches of annotation acquisition, can be summarized as four types in general, i.e., marker-based [26, 27, 28, 29], synthetic [30, 31, 32, 33, 34], manual [30, 35] or hybrid [36, 37, 38, 39, 40], and automatic [41, 42, 43, 44] annotated datasets.
-Marker-based approaches, using magnetic sensor [26, 27] or Mocap markers [28, 29], can alter and induce bias to the hand appearance.
-Synthetic data [30, 31, 33, 34, 32] suffers from the sim2real gap in terms of hand motion and texture features.
-Introducing human annotators to fully [30, 35] or partly [36, 37, 38, 39, 40] annotate 2D/3D keypoints circumvents the issues above, but it either limits the scale of datasets or manifests costly and laborious in practice.
-Most similar to ours, some datasets adopt fully automatic pipelines to obtain 3D hand pose annotations [41, 42, 43], which leverage pre-trained models (e.g.
-OpenPose [45]) to infer the prior hand information and rely on optimization to fit the MANO hand model [25].
-Despite the existing progress, previous datasets only provide depth [26, 32], RGB images [33, 38, 37, 40] or both of them [27, 28, 31, 34, 30, 35, 36, 39, 41, 42, 43] as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation.
+Existing datasets, according to their approaches of annotation acquisition, can be summarized as four types in general, i.e., marker-based (yuan2017bighand2, ; garcia2018first, ; taheri2020grab, ; fan2023arctic, ), synthetic (mueller2017real, ; hasson2019learning, ; mueller2019real, ; mueller2018ganerated, ; zimmermann2017learning, ), manual (mueller2017real, ; sridhar2016real, ) or hybrid (chao2021dexycb, ; zimmermann2019freihand, ; moon2020interhand2, ; liu2022hoi4d, ; ohkawa2023assemblyhands, ), and automatic (hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ; sener2022assembly101, ) annotated datasets.
+Marker-based approaches, using magnetic sensor (yuan2017bighand2, ; garcia2018first, ) or Mocap markers (taheri2020grab, ; fan2023arctic, ), can alter and induce bias to the hand appearance.
+Synthetic data (mueller2017real, ; hasson2019learning, ; mueller2018ganerated, ; zimmermann2017learning, ; mueller2019real, ) suffers from the sim2real gap in terms of hand motion and texture features.
+Introducing human annotators to fully (mueller2017real, ; sridhar2016real, ) or partly (chao2021dexycb, ; zimmermann2019freihand, ; moon2020interhand2, ; liu2022hoi4d, ; ohkawa2023assemblyhands, ) annotate 2D/3D keypoints circumvents the issues above, but it either limits the scale of datasets or manifests costly and laborious in practice.
+Most similar to ours, some datasets adopt fully automatic pipelines to obtain 3D hand pose annotations (hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ), which leverage pre-trained models (e.g.
+OpenPose (8765346, )) to infer the prior hand information and rely on optimization to fit the MANO hand model (MANO, ).
+Despite the existing progress, previous datasets only provide depth (yuan2017bighand2, ; mueller2019real, ), RGB images (mueller2018ganerated, ; moon2020interhand2, ; zimmermann2019freihand, ; ohkawa2023assemblyhands, ) or both of them (garcia2018first, ; taheri2020grab, ; hasson2019learning, ; zimmermann2017learning, ; mueller2017real, ; sridhar2016real, ; chao2021dexycb, ; liu2022hoi4d, ; hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ) as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation.
 ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images.
 Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications.
-Figure 1: Data capture setup with the HMSP and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
-2.2 Image-based 3D Hand Pose Estimation
-As a key computer vision task, 3D hand pose estimation from images is highly demanded by applications like XR [1, 2, 3, 4], human-robot interaction [5, 6, 7] and imitation learning [8, 9, 10].
-Therefore, this field has been extensively explored in previous arts that uses single RGB [46, 47, 48, 49, 50, 13, 38, 14, 51, 52, 53, 31, 54, 55, 56, 57] or depth [58, 59, 60, 61, 62] image as input.
+2.2. Image-based 3D Hand Pose Estimation
+As a key computer vision task, 3D hand pose estimation from images is highly demanded by applications like XR (wu2020hand, ; sagayam2017hand, ; marchand2015pose, ; liang2015ar, ), human-robot interaction (sampieri2022pose, ; gao2021dynamic, ; gao2019dual, ) and imitation learning (qin2022dexmv, ; chen2023bi, ; antotsiou2018task, ).
+Therefore, this field has been extensively explored in previous arts that uses single RGB (jiang2023a2j, ; spurr2020weakly, ; yang2019disentangling, ; cheng2021handfoldingnet, ; kim2021end, ; lin2021two, ; moon2020interhand2, ; li2022interacting, ; smith2020constraining, ; wang2020rgb2hands, ; zhang2021interacting, ; hasson2019learning, ; liu2021semi, ; boukhayma20193d, ; baek2019pushing, ; zhang2019end, ) or depth (ge2018robust, ; moon2018v2v, ; xiong2019a2j, ; zhang2020handaugment, ; zhu2019hand, ) image as input.
 These methods can be roughly categorized into two fashions, i.e., model-based and model-free methods.
-Model-based methods [14, 51, 52, 53, 31, 54, 55, 56, 57] utilize the prior knowledge of the MANO hand model [25] by estimating its shape and pose parameters, while model-free methods [46, 47, 48, 49, 50, 13, 38, 58, 59, 60, 61, 62] learn the direct regression of 3D hand joints or vertices coordinates.
-Recently there has been growing interest in leveraging the temporal supervision [63, 54, 64, 65, 66] or leveraging sequential images as input [15, 16, 17, 67, 68, 69] for 3D hand pose estimation.
+Model-based methods (li2022interacting, ; smith2020constraining, ; wang2020rgb2hands, ; zhang2021interacting, ; hasson2019learning, ; liu2021semi, ; boukhayma20193d, ; baek2019pushing, ; zhang2019end, ) utilize the prior knowledge of the MANO hand model (MANO, ) by estimating its shape and pose parameters, while model-free methods (jiang2023a2j, ; spurr2020weakly, ; yang2019disentangling, ; cheng2021handfoldingnet, ; kim2021end, ; lin2021two, ; moon2020interhand2, ; ge2018robust, ; moon2018v2v, ; xiong2019a2j, ; zhang2020handaugment, ; zhu2019hand, ) learn the direct regression of 3D hand joints or vertices coordinates.
+Recently there has been growing interest in leveraging the temporal supervision (hasson2020leveraging, ; liu2021semi, ; yang2020seqhand, ; Kocabas_2020_CVPR, ; park2022handoccnet, ) or leveraging sequential images as input (wen2023hierarchical, ; Fu_2023_ICCV, ; cho2023transformer, ; Cai_2019_ICCV, ; khaleghi2022multi, ; choi2021beyond, ) for 3D hand pose estimation.
 In this study, we evaluate existing methods and our baseline method in both single image-based and video-based problem settings, respectively.
 Apart from the previous approaches, we investigate the potential of thermal imagery for tackling various challenges in 3D hand pose estimation.
-2.3 Thermal Computer Vision
-Thermal cameras achieve imaging by capturing the radiation emitted in the LWIR spectrum and deducing the temperature distribution on the surfaces [24].
-Leveraging its robustness to variable illumination and unique temperature information, numerous efforts have been made to address various computer vision tasks, including super-resolution [70, 71, 72], human detection [73, 74, 75], action recognition [76, 77] and pose estimation [78, 79, 80], semantic segmentation [81, 82, 83, 84], depth estimation [85, 86, 87, 88], visual(-inertial) odometry/SLAM [89, 90, 91, 92], 3D reconstruction [93, 94, 95], etc.
+Figure 2. Design of the head-mounted sensor platform and sensor alignment.
+2.3. Thermal Computer Vision
+Thermal cameras achieve imaging by capturing the radiation emitted in the LWIR spectrum and deducing the temperature distribution on the surfaces (lloyd2013thermal, ).
+Leveraging its robustness to variable illumination and unique temperature information, numerous efforts have been made to address various computer vision tasks, including super-resolution (rivadeneira2023thermal, ; gupta2021toward, ; kansal2020multi, ), human detection (10.5555/3451271.3451272, ; ivavsic2019human, ; haider2021human, ), action recognition (batchuluun2019action, ; ding2022individual, ) and pose estimation (lupion20243d, ; chen2020multi, ; smith2023human, ), semantic segmentation (kutuk2022semantic, ; kim2021ms, ; vertens2020heatnet, ; li2020segmenting, ), depth estimation (Shin_2023_CVPR, ; lu2021alternative, ; shin2021self, ; kim2018multispectral, ), visual(-inertial) odometry/SLAM (rob.21932, ; 9623261, ; shin2019sparse, ; saputra2020deeptio, ), 3D reconstruction (liu2023humans, ; sage20213d, ; schramm2022combining, ), etc.
 In this work, we focus on 3D hand pose estimation, which is an under-exploited task based on thermal images.
-3 The ThermoHands Benchmark
-3.1 Multi-Spectral Hand Pose Dataset
+Figure 3.
+Thermal calibration chessboard containing a black base board and multiple removable white cubes (a).
+By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
+3. The ThermoHands Benchmark
+3.1. Multi-Spectral Hand Pose Dataset
 Overview.
 At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab.
-1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from School of Informatics, University of Edinburgh, and participant consent forms were signed before the collection..
+1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^3] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection..
 As shown in Fig.
@@
 To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
-Normal office (Main) Other settings Total
-Setting train val test sum darkness sun glare gloves kitchen
-#frames 47,436 12,914 24,002 84,352 3,188 2,508 3,068 2,808 95,924
-#seqs 172 43 86 301 12 12 12 14 352
-#subjects 16 4 8 28 1 1 1 2 -
-Table 1: Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
-Figure 2: Thermal calibration chessboard containing a black base board and multiple removable white cubes (a).
-By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
 Sensor Platforms.
-The HMSP is mounted with an Intel RealSense L515 LiDAR depth camera [20] streaming egocentric RGB, depth and NIR images, and a FLIR Boson 640 thermal camera [96] receiving the LWIR to obtain the thermal images.
-To better simulate XR devices, the HMSP is modularly designed as two components: a base and a sensor board with a fixed 30-degree downward tilt.
-An extra exocentric platform equipped with an Intel RealSense D455 [21] is leveraged to support multi-view annotation (cf.
+Apart from the traditional approach of mounting cameras on helmets (kwon2021h2o, ), the design of the HMSP (cfFig.
+2) focuses on simulating an actual Mixed Reality (MR) device and reducing extra weight to allow participants to perform freely.
+The HMSP consists of three major components: a cushion for comfort, a base component that provides a 30-degree downward tilt, and a sensor board that carries two cameras - an Intel RealSense L515 LiDAR camera (IntelRSL515, ) that streams egocentric RGB, depth, and NIR images, and a Teledyne FLIR Boson 640 long-wave infrared (LWIR) camera (FLIRBoson2024, ) that receives the LWIR to obtain the thermal images.
+An extra exocentric platform equipped with an Intel RealSense D455 (IntelRealSense2023D455, ) is leveraged to support multi-view annotation (cf.
 Sec.
-3.2) as well as provide the RGB-D image data from the third-person viewpoint.
+4) as well as provide the RGB-D image data from the third-person viewpoint.
 As exhibited in Fig.
-1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters [20, 21].
-In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
+1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters (IntelRSL515, ; IntelRealSense2023D455, ).
 Synchronization.
@@
 the timestamps of thermal images (8.5fps), thereby generating well-aligned multi-spectral, multi-view data samples as our released data.
-Calibration.
-For accuracy, we use factory-calibrated intrinsic parameters from the D455 and L515 cameras [20, 21].
-To better calibrate the parameters of the thermal camera, we self-design a modular calibration chessboard as shown in Fig.
-2.
+Egocentric Calibration.
+For accuracy, factory-calibrated intrinsic parameters are used for the Intel RealSense D455 and L515 cameras (IntelRSL515, ; IntelRealSense2023D455, ).
+As seen in Fig.
+2, although the thermal camera’s center is aligned with the RGB camera’s center at the same height on the 3D-printed sensor board, we still calibrate the extrinsic parameters between these two egocentric cameras to ensure enhanced alignment accuracy.
+To accommodate the unique imaging mechanism of thermal camera, we self-design a modular calibration chessboard as shown in Fig.
+3.
 Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig.
-2.
-In this way, we can simultaneously calibrate the intrinsic parameter of the L515 thermal camera and its extrinsic parameter w.r.t.
+3.
+In this way, we can simultaneously calibrate the intrinsic parameter of the thermal camera and its extrinsic parameter w.r.t.
 the D455 RGB-D camera.
-To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table.
-At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP problem [97].
-For the subsequent frames, as the exocentric platform keeps stationary, we only track the the motion of the ego-head using KISS-ICP [98] odometry method, which takes the point clouds converted from the egocentric depth image as input.
+Cross-View Calibration.
+To calibrate the cameras between the egocentric and exocentric viewpoints, we place 11 IR reflective spheres at random heights within the playground, ensuring they are visible from both viewpoints.
+Initially, we plan to detect these spheres automatically from the NIR images and track them with the Kalman Filter (wan2000unscented, ).
+However, we find this approach leads to many false positive detections, which severely affect the tracking accuracy.
+To ensure the calibration accuracy, we only annotate the sphere markers manually from two viewpoints in the first frame.
+The 3D positions of markers can be computed using the egocentric depth image and the cross-view transformation can be computed by solving the PnP (FischlerBolles1981, ).
+For the subsequent frames, we can calculate the transformation by only tracking the pose of the egocentric camera as the exocentric platform keeps stationary during collection.
+We leverage the state-of-the-art odometry method KISS-ICP (vizzo2023kiss, ) and track the motion of the egocentric camera with the point clouds converted from depth images.
+Specifically, we set the initial_threshold and min_motion_th parameters of KISS-ICP as 0.0001 to make it capable of catching subtle motion.
+The input point cloud range is set as [0.2, 2.0] meters while the max_points_per_voxel is 30.
 Dataset Statistics.
@@
 We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen.
-Please refer to the supplementary for more dataset details.
-3.2 Hand Pose Annotation
-To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [41, 42, 43], to obtain the 3D hand pose ground truth for our dataset.
-In particular, we use the MANO statistical hand model [25] to represent 3D hand pose.
-The MANO model parameterizes the hand mesh vertices into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$.
+Figure 4.
+Automatic annotation pipeline of 3D hand pose.
+We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands (mediapipe_hands_2020, ) and SAM (kirillov2023segment, ).
+Various error terms are formulated to optimize the MANO parameters.
+4. Hand Pose Annotation
+To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in (hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ), to obtain the 3D hand pose ground truth for our dataset.
+In particular, we use the MANO statistical hand model (MANO, ) to represent 3D hand pose.
+The MANO model parameterizes the hand mesh vertices $\mathcal{V}(\beta,\theta)$ into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$, consisting of 45 parameters accounting for 15 hand joint angles (3 DoF for each) plus the rest for global rotation and translation.
+Following the MANO PyTorch version in (hasson19_obman, ), we denote the $N_{\mathcal{J}}=21$ hand joints mapped from the hand parameters as $\mathcal{J}(\beta,\theta)$.
 The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
-$\theta^{*}=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$ (1)
-where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors.
-The diagram illustrating our annotation process is shown in Fig.
-3.
-For more explanation of error terms and implementation details of our dataset annotation process, please refer to our supplementary document.
-Figure 3: Automatic annotation pipeline of 3D hand pose.
-We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands [99] and SAM [100].
-Various error terms are formulated to optimize the MANO parameters.
-3.3 TherFormer: A Baseline Method
+(1) $\theta^{*}$ $=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$
+$+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$
+where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 4.
+Initialization.
+For each sequence, we optimize the shape parameter $\beta$ [^5] For simplicity, we omit $\beta$ in Eq.
+1 and subsequent equations.
+together with $\theta$ only at the first frame using Eq.
+1 until convergence and fix its values for the subsequent frames.
+We initialize the pose parameter $\theta$ for each frame using the optimization result from the last frame.
+This helps to accelerate the convergence as well as keep the temporal consistency.
+2D Joint Error $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}$.
+Given RGB images from $N_{C}$ viewpoints, we infer 2D hand joints $\mathcal{J}^{2D}$ with MediaPipe Hands (mediapipe_hands_2020, ) and define the 2D joint error as:
+(2) $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{2D}_{c,i}-\pi_{c}(\mathcal{J}(\theta)_{i})||$
+where $\pi_{c}(\cdot)$ returns the 2D projection location for 3D position in the $c$-th camera viewpoint, and $\alpha_{c}$ is hyperparameter used to weigh different viewpoints.
+Figure 5.
+Overall Framework of TherFormer.
+Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction.
+Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
+2D Mask Error $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}$.
+To generate the high-quality 2D hand mask, we prompt the prevalent Segment Anything Model (kirillov2023segment, ) with the 2D hand joints $\mathcal{J}^{2D}$ and the bounding box derived from it.
+We penalize the distance between the hand mesh vertices $\mathcal{V}(\theta)$ and the binary hand mask $\mathcal{M}_{c}$ as:
+(3) $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{M}_{c,j}-\pi_{c}(\mathcal{V}(\theta)_{i})||$
+where $\mathcal{M}_{c,j}$ is the coordinate of $j$-th non-zero pixel in the mask $\mathcal{M}_{c}$.
+3D Joint Error $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$.
+We triangulate the 2D joints from multiple views to lift them to 3D joints $\mathcal{J}^{3D}$ and measure their difference to $\mathcal{J}(\theta)$, which is written as:
+(4) $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}=\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{3D}_{i}-\mathcal{J}(\theta)_{i}||$
+3D Mesh Error $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}$.
+To better supervise the hand mesh and fasten the optimization, we generate the 3D hand pose cloud $\mathcal{P}$ by cropping the depth image using the 2D hand mask and merging all views together.
+The 3D mesh error term compensates for the distance between the hand mesh and point cloud:
+(5) $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}=\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{P}_{j}-\mathcal{V}(\theta)_{i}||$
+Regularization $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$.
+To alleviate irregular hand articulation, we constrain the joint angles to pre-defined lower and upper boundaries $\underline{\theta}$ and $\overline{\theta}$:
+(6) $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}=\sum_{i=1}^{45}(\max(\underline{\theta_{i}}-\theta_{i},0)+\max(\theta_{i}-\overline{\theta_{i}},0))$
+Note that we also impose regularization to the shape parameter $\beta$ during initialization. For more implementation details of our dataset annotation process, please refer to Appendix B.
+5. TherFormer: A Baseline Method
 For our benchmark, we setup a baseline method, dubbed TherFormer, for thermal image-based 3D hand pose estimation.
 As exhibited in Fig.
-4, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient.
+5, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient.
 Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images.
 However, our primary objective is to establish a baseline method for future research, rather than focusing exclusively on methods for thermal images.
-Figure 4: Overall Framework of TherFormer.
-Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction.
-Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
 Problem Definition.
@@
 Given backbone features, we first utilize a mask head to estimate the binary hand mask in the thermal image.
-Then, we leverage the deformable self-attention [101] to refine the hand spatial features under the guidance of the estimated hand mask.
+Then, we leverage the deformable self-attention (zhu2020deformable, ) to refine the hand spatial features under the guidance of the estimated hand mask.
 Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations.
@@
 Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities.
-To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention [102] to explicitly attend to the feature vector of every frame.
+To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention (vaswani2017attention, ) to explicitly attend to the feature vector of every frame.
 The output is frame-wise spatio-temporal feature embeddings.
@@
 We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cfSec.
-3.2).
+4).
 For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately.
-Please see more details of TherFormer in our supplementary document.
-4 Experiment
-4.1 Evaluation of the Annotation Method
+Ego-view optimization Multi-view optimization
+Errors $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}+\mathcal{L}_{j3d}$
+mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
+Table 2. Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
+Figure 6. Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
+     Method      Input      MEPE (mm) $\downarrow$      AUC $\uparrow$      MEPE-RA (mm) $\downarrow$      AUC-RA $\uparrow$      fps $\uparrow$
+     (a)      A2J-Transformer (jiang2023a2j, )      Single      51.68      0.474      20.76      0.603      34
+     (b)      HTT (wen2023hierarchical, )      Single      49.09      0.489      20.69      0.599      211
+     (c)      TherFormer-S      Single      44.64      0.539      18.34      0.643      136
+     (d)      (c) w/o spatial transformer      Single      48.79      0.491      20.15      0.609      174
+     (e)      (c) w/o mask guidance      Single      48.83      0.494      18.89      0.625      141
+     (f)      HTT (wen2023hierarchical, )      Sequence      47.07      0.512      17.49      0.659      129
+     (g)      TherFormer-V      Sequence      43.36      0.549      17.36      0.661      52
+Table 3. Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
+6. Experiment
+6.1. Evaluation of the Annotation Method
 As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec.
-3.2) and analyze the impact for optimization results.
+4) and analyze the impact for optimization results.
 For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames.
-To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D joint positions by triangulation.
+To that end, we first annotate the 2D joint locations on RGB images from two viewpoints and obtain the 3D joint positions by triangulation.
 We calculate the average 3D joint errors across all frames to measure the accuracy.
 As shown in Tab.
-2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [43, 41, 37].
+2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of (kwon2021h2o, ; hampali2020honnotate, ; zimmermann2019freihand, ).
 The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture.
@@
 We showcase some annotation examples in Fig.
-5.
+6.
 As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subjects’ hand color and shape.
-Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
-Ego-view optimization Multi-view optimization
-Errors $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}+\mathcal{L}_{j3d}$
-mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
-Table 2: Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
-4.2 Experiment Setup
+Please see our project page for more visualization of annotation.
+Figure 7.
+Impact of the temporal sequence length $T$ to TherFormer.
+The plots show all four metrics MEPE (mm) $\downarrow$, AUC $\uparrow$, MEPE-RA (mm) $\downarrow$, and AUC-RA $\uparrow$ against 5 sequence length settings, i.e., $T\in${1, 2, 4, 8, 16}.
+6.2. Experiment Setup
 Dataset Preparation.
@@
 1) automatically following Sec.
-3.2, of which the training and validation sets serve as the foundation for the training of all network models.
+4, of which the training and validation sets serve as the foundation for the training of all network models.
+Our automatic annotation pipeline (cf.
+Sec.
+4) becomes infeasible under challenging scenarios, i.e., gloves, darkness and sun glare, since hands appear corrupted in either RGB or depth images.
+To facilitate the quantitative evaluation under challenging settings, we manually annotate the ground truth for a few sequences collected in the glove and sun glare scenarios (cfTab.
+1).
+Specifically, we first annotate the 2D keypoints from two viewpoints and then use triangulation to obtain their 3D positions.
 Method and Implementation.
-To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT [15] and A2J-Transformer [46], and reproduced them on our dataset for experiments.
-HTT [15] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [46] only works under the single image-based setting.
+To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT (wen2023hierarchical, ) and A2J-Transformer (jiang2023a2j, ), and reproduced them on our dataset for experiments.
+HTT (wen2023hierarchical, ) is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer (jiang2023a2j, ) only works under the single image-based setting.
 For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines.
-We exclude the additional action block used by HTT [15], focusing solely on hand pose estimation.
-We adjusted the anchor initialization phase of A2J-Transformer [46] to better accommodate our dataset, without altering the density of its anchors.
+We exclude the additional action block used by HTT (wen2023hierarchical, ), focusing solely on hand pose estimation.
+We adjusted the anchor initialization phase of A2J-Transformer (jiang2023a2j, ) to better accommodate our dataset, without altering the density of its anchors.
 All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
 Evaluation Metrics.
-We evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) [34], in both camera space and root-aligned (RA) space [38].
-For RA space, we align the estimated wrist with its groundtruth position before measurement.
-For PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
-4.3 Thermal Image-based 3D Hand Pose Estimation
+Following HTT (wen2023hierarchical, ), we evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) (zimmermann2017learning, ), in both camera space and root-aligned (RA) space.
+For the RA space, we align the estimated wrist with its groundtruth position before measuring the joint errors.
+In terms of PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
+6.3. Thermal Image-based Results
 Main Results.
-To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods [15, 46] on the main testing set, as shown in Tab.
+To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods (wen2023hierarchical, ; jiang2023a2j, ) on the main testing set, as shown in Tab.
 3.
-TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT [15] given the same sequential images as input.
+TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT (wen2023hierarchical, ) given the same sequential images as input.
 Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec.
-3.3).
-A performance gap can be observed between single image-based and video-based settings for both HTT [15] and TherFormer.
+5).
+A performance gap can be observed between single image-based and video-based settings for both HTT (wen2023hierarchical, ) and TherFormer.
 We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities.
 Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices.
-Moreover, TherFormer-V also improves the performance over HTT [15] for other spectra (cfTab.
-4), proving its adaptability to different inputs.
-Method Input MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$ fps $\uparrow$
-(a) A2J-Transformer [46] Single 51.68 0.474 20.76 0.603 34
-(b) HTT [15] Single 49.09 0.489 20.69 0.599 211
-(c) TherFormer-S Single 44.64 0.539 18.34 0.643 136
-(d) (c) w/o spatial transformer Single 48.79 0.491 20.15 0.609 174
-(e) (c) w/o mask guidance Single 48.83 0.494 18.89 0.625 141
-(f) HTT [15] Sequence 47.07 0.512 17.49 0.659 129
-(g) TherFormer-V Sequence 43.36 0.549 17.36 0.661 52
-Table 3: Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
+Moreover, TherFormer-V also improves the performance over HTT (wen2023hierarchical, ) for other spectra (cfTab.
+5), proving its adaptability to different inputs.
+Figure 8. Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
+TherFormer-V (glove) TherFormer-V (sun glare)
+Spectrum MEPE-RA (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC $\uparrow$
+RGB 51.94 0.141 38.24 0.252
+Depth 45.96 0.206 42.27 0.254
+NIR 39.83 0.282 90.84 0.093
+Thermal 39.23 0.302 32.56 0.363
+Table 4. Comparison between different spectra under two challenging conditions, i.e., glove and sun glare.
+HTT (Sequence) TherFormer-V TherFormer-S Best
+Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$
+RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.30 0.542
+Depth 41.62 0.559 39.70 0.581 39.84 0.579 39.70 0.581
+NIR 41.57 0.562 40.79 0.575 40.98 0.575 40.79 0.575
+Thermal 47.07 0.512 43.36 0.549 44.64 0.539 43.36 0.549
+Table 5. Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
 Ablation Studies.
-We ablate our proposed mask-guided attention mechanism and the entire spatial transformer (cfSec.
-3.3) to observe their impact.
+We ablate our proposed mask-guided attention mechanism and the entire spatial transformer to observe their impact.
 As can be seen in Tab.
@@
 (d)), confirming the importance of spatial feature enhancement for fine-grained hand joints localization.
-4.4 Comparison between Spectrum
-4.4.1 Qualitative Results under Challenging Conditions
-To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf.
+Impact of Sequence Length.
+Here we conduct experiments to analyze the impact of the sequence length $T$ on the performance of our baseline method.
+We present the variation in our performance with respect to sequence length in Fig.
+7.
+All experiments are conducted using the same hyperparameters, except for the $T=16$ model, which uses a smaller batch size due to GPU memory constraints.
+It can be observed that performance generally improves with increased sequence length, particularly in root-aligned metrics.
+There is a slight performance decrease in the camera space when the sequence length $T$ is increased to 16 but the performance in the root-aligned space is improved.
+The general trend underscores the importance of temporal information in enhancing the accuracy of 3D hand pose estimation.
+6.4. Comparison between Spectrum
+6.4.1. Qualitative Results under Challenging Conditions
+To justify the advantages of using thermal images for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on sequences collected in challenging scenarios (cf.
 Tab.
 1).
-Our automatic annotation pipeline (cf.
-Sec.
-3.2) becomes infeasible since hands appear corrupted in either RGB or depth images.
-Therefore, we conduct a qualitative analysis of their performance and present some representative examples in Fig.
-6.
-We use the same model, i.e., HTT [15], trained on different spectra for a fair comparison.
-Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
+To intuitively show the results, we first conduct a qualitative analysis of their performance and present some representative examples in Fig.
+8.
+We use the same model, i.e., HTT (wen2023hierarchical, ), trained on different spectra for a fair comparison.
+Please refer to our project page for more figure examples and demo videos under challenging conditions.
 RGB vs.
@@
 Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process.
-Contrary to RGB techniques, as shown in Fig.
+Contrarily, as shown in Fig.
 8 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
@@
 The thermal camera’s ability to consistently capture hand features in diverse lighting conditions positions it as a suitable option for future XR applications.
-Figure 6: Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
-4.4.2 Quantitative Results under Normal Conditions
+6.4.2. Quantitative Results under Challenging Conditions
+To provide the numerical results, we evaluate our TherFormer-V models trained for different spectra on our manually annotated sequences (cfSec.
+6.2) and calculated the quantitative results presented in Tab.
+4.
+As can be seen, thermal imaging-based approaches demonstrate the best performance among different spectra in challenging settings, underscoring thermal imagery’s advantages in difficult lighting conditions and when hands are occluded.
+Notably, the conclusion drawn from Tab.
+4 resonate with the visualization in Fig.
+8.
+The RGB spectrum performs poorly in scenarios involving gloves, where the color and texture of human hands are significantly altered.
+Intense sunlight introduces severe artifacts in NIR images, markedly impairing performance under sun glare.
+6.4.3. Quantitative Results under Normal Conditions
 To validate the versatility of thermal-based methods, we also evaluate their performance compared to other spectra (i.e., RGB, depth, NIR and thermal) under normal conditions for general-purpose use cases.
 We apply different spectral images as input to multiple baselines methods and report the results in Tab.
-4.
+5.
 Not surprisingly, depth image-based methods yield the best performance on average, as they directly utilize the depth information to provide a detailed 3D structure of the hand.
@@
 However, active NIR sensors are prone to interference from external NIR sources like sun glare (cfFig.
-6).
+8).
 Thermal-based methods, despite using single-channel heat capture, achieves comparable performance to those using RGB images that contain rich color and textures.
 Such results suggest that thermal images can not only serve as supplements in challenging cases but also can be a viable alternative to other spectra in normal conditions for hand pose estimation.
-HTT (Sequence) TherFormer-V TherFormer-S Average
-Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$
-RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.80 0.538
-Depth 41.62 0.559 39.70 0.581 39.84 0.579 40.39 0.573
-NIR 41.57 0.562 40.79 0.575 40.98 0.575 41.11 0.571
-Thermal 47.07 0.512 43.36 0.549 44.64 0.539 45.06 0.533
-Table 4: Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
-5 Limitation and Future Work
-Despite the advancements, we acknowledge certain limitations here guiding our future work.
-First, the scale of our current dataset is relatively small when compared to some datasets [29, 40, 39, 38].
-We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions.
-Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands.
-Further efforts will be annotating the fine-grained hand action splits and contact heatmap for hand grasp [42].
-Lastly, we currently rely on qualitative analysis for challenging conditions when our automatic annotation method becomes unreliable.
-In the future, we will leverage unaffected spectral images in individual cases to manually annotate the 3D hand pose.
-6 Conclusion
+7. Conclusion
 This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images.
-It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships.
+ThermoHands features a multi-spectral, multi-view dataset with automatically annotated 3D hand poses and a novel baseline method, TherFormer, utilizing dual transformer modules for encoding spatio-temporal relationships.
 We demonstrate near 1cm annotation accuracy, show that TherFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios.
-We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.
-Acknowledgments and Disclosure of Funding
-This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1), and partially supported by Cisco Research.
-Checklist
-  1. 1.
-    For all authors…
-    1. (a)
-      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-[Yes] We elaborate our contributions as a list in the introduction section and briefly explain them in the abstract.
-    2. (b)
-      Did you describe the limitations of your work? [Yes] See Sec. 5
-    3. (c)
-      Did you discuss any potential negative societal impacts of your work? [N/A] There is no societal impact of this work.
-    4. (d)
-      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We have reviewed the ethics review guidelines and make sure we obey all rules for this research.
-  2. 2.
-    If you are including theoretical results…
-    1. (a)
-      Did you state the full set of assumptions of all theoretical results? [N/A] This work does not contribute to the theoretical results.
-    2. (b)
-      Did you include complete proofs of all theoretical results? [N/A] This work does not contribute to the theoretical results.
-  3. 3.
-    If you ran experiments (e.g. for benchmarks)…
-    1. (a)
-      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
-[No] We will release our code upon the acceptance of this paper.
-    2. (b)
-      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We describe them in our main paper and supplementary document.
-    3. (c)
-      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
-[No] In line with the state-of-the-art 3D hand pose estimation works, we do not report error bars for our experiment results.
-    4. (d)
-      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We include the type of GPUs in the main paper.
-  4. 4.
-    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
-    1. (a)
-      If your work uses existing assets, did you cite the creators? [Yes] We use the offical code of HTT [15] and A2J-transformer [46] to reproduce their results on our dataset, which have been cited.
-    2. (b)
-      Did you mention the license of the assets? [Yes] We mention the license of the offical codes of HTT and A2J-Transformer in the supplementary document.
-    3. (c)
-      Did you include any new assets either in the supplemental material or as a URL? [Yes] We include the datasheet and the download URL of our dataset in the supplemental document.
-    4. (d)
-      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
-[Yes] We claim in the footnote that participant consent forms were signed before the collection.
-    5. (e)
-      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We discuss in the supplementary document.
-  5. 5.
-    If you used crowdsourcing or conducted research with human subjects…
-    1. (a)
-      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We do not give written instructions to our participants.
-    2. (b)
-      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] There is no potential participant risks.
-    3. (c)
-      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] Our participants are volunteer.
+We believe our foundational endeavour could set the stage for further research in thermal-based 3D hand pose estimation and its wide application.
+Limitation.
+As the first exploration in this field, this work has certain limitations that can guide our future work.
+For example, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands.
+Further efforts could be annotating the fine-grained hand action splits and hand-object contact (brahmbhatt2020contactpose, ).
+Appendix
+The appendix is organized as follows:
+  • $\bullet$
+    Appendix A illustrates more details about our hand pose dataset, in the aspects of spectrum coverage, interference avoidance, and dataset distribution.
+  • $\bullet$
+    Appendix B presents more implementation details of our automatic 3D hand pose annotation method.
+  • $\bullet$
+    Appendix C introduce more architecture design, parameter and loss details of TherFormer.
+We also provide more visualization in the format of both images and videos for our data, annotation and results on our anonymous project page.
+Appendix A Multi-spectral Hand Pose Dataset
+Spectrum Coverage and Interference Avoidance.
+The specification of all sensor frames we collect are shown in Tab.
+1.
+Note that we keep the faces of subjects outside of the FoV of the exocentric camera to avoid any personally identifiable information.
+The multi-spectral data encompasses imaging from the visible spectrum (400-700nm), near-infrared (NIR) spectrum (850 nm $\pm$ 10 nm), and the LWIR spectrum (8-14 $\mu$m).
+Images captured across different spectra contain unique information and serve varied purposes.
+For instance, RGB images offer semantic information, facilitating a deeper understanding of human-environment interaction.
+NIR lights can be actively emitted by our depth cameras to obtain the depth measurements via ToF or structured lighting.
+The LWIR frame, capturing temperature information, readily isolates uniform heat emitters like human hands.
+On the head-mounted sensor platform, the L515 LiDAR depth camera (IntelRSL515, ) emits NIR lasers at a wavelength of 860 nm, which falls outside the thermal camera’s receptive range (8-14 $\mu$m), thereby eliminating any potential interference between cameras on the HMSP.
+Conversely, the exocentric RGB-D camera (IntelRealSense2023D455, ) necessitates structured lighting employing NIR at a wavelength identical to the IR emitter on the L515 LiDAR camera (IntelRSL515, ).
+To prevent interference and image corruption, the exocentric RGB-D camera and the egocentric NIR LiDAR are strategically positioned outside each other’s receptive fields during data collection.
+In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
+Resolution Fov FPS
+Sensor frames Sensor Type Range Horizontal Vertical Range Horizontal Vertical
+RGB (ego) Intel RS L515 (IntelRSL515, ) - 1280 720 - 69 42 30
+NIR (ego) Intel RS L515 (IntelRSL515, ) - 640 480 - 70 55 30
+Depth (ego) Intel RS L515 (IntelRSL515, ) ¡ 5mm @ 1m 640 480 0.25m to 9m 70 55 30
+Thermal (ego) FLIR Boson 640 (FLIRBoson2024, ) $\leq$ 60 mK 640 512 - 95 - 8.5
+RGB (exo) Intel RS D455 (IntelRealSense2023D455, ) - 1280 720 - 87 58 30
+[…]
+Table 6. The specification of sensor frames captured in data collection.
+Figure 9. Distribution of data from the main part over hand actions.
+Dataset Distribution.
+We show the data distribution of our main part over hand actions in Fig.
+9.
+Among hand-object interaction actions, write with pen, write with pencil and read book have more frames than others due to the complexity of these actions, making them tend to last longer than others.
+As seven of our participants did not perform the hand-virtual actions, we have fewer frames from them than their hand-object counterparts.
+Please see our project page for more visualization of our collected data.
+Appendix B Dataset Annotation
+Acquisition of 2D Keypoint and Mask.
+We utilize the open-source MediaPipe Hands pipeline (mediapipe_hands_2020, ) to infer the 2D hand keypoints from both egocentric and exocentric RGB images.
+To improve the recall of hand detection, we modify the hyperparameters min_detection_confidence from the default 0.5 to 0.1.
+Sequential RGB images are fed into the MeidaPipe Hands pipeline to obtain the 2D keypoints for two hands.
+Particularly, we distinguish between two hands according to their relative locations on images.
+Given 2D keypoints estimated from the task-specific MeidiaPipe Hands model, we employ the versatile Segment-Anything Model (SAM) (li2020segmenting, ) to infer the 2D hand masks.
+To ensure the high quality of the generated 2D hand mask, we utilize the largest version of SAM, i.e., ViT-L SAM model and prompt it with both the 2D hand keypoints and the bounding boxes defined by them.
+As a result, we acquire 2D hand keypoints and masks of two hands for each frame.
+3D Keypoint Triangulation.
+After inferring the 2D hand keypoints from two views, we can obtain the positions of 3D hand keypoints using triangulation.
+To implement this, we use the OpenCV function cv2.triangulatePoints (opencv_library, ).
+3D Hand Point Cloud Generation.
+To generate 3D hand point cloud, we first index the hand pixels with the 2D hand mask on the depth image and then convert them into 3D points with the camera intrinsic parameters.
+Points generated from the exocentric view are transformed into the egocentric camera space.
+By merging points from two viewpoints, we can obtain a dense 3D hand point cloud.
+Joint Angle Limitation.
+Following (hampali2020honnotate, ; kwon2021h2o, ; liu2022hoi4d, ), we optimize the MANO model in the joint angle space instead of the PCA space so that we can constrain the joint angles explicitly.
+As said in the main paper, the joint angles are limited by a set of empirical boundary values during optimization.
+Specifically, we use the same parameters of the upper and lower joint angle boundaries as (kwon2021h2o, ).
+Shape Regularization. To avoid unrealistic hand shape, we also add regularization to the shape parameter $\beta$ when optimizing it for the first frame of each sequence, which can be written as:
+(7) $\mathcal{L}_{shape}=\sum_{i=1}^{10}||\beta_{i}||$
+MANO Fitting.
+We utilize the Pytorch version of the MANO layer proposed in (hasson19_obman, ).
+The Adam optimizer (kingma2014adam, ) is used to minimize the optimization objective mentioned in the main paper.
+To obtain accurate 3D hand pose annotation, we optimize the MANO parameters individually for each frame instead of considering batches of them.
+For the first frame of each sequence, we run the optimization for 500 iterations with an initial learning rate of as 0.1 decayed by 0.9 for each 50 frames.
+For the subsequent frames, as we initialize with the results from the last frame, we only optimize for 60 iterations and initialize the learning rate as 0.05 to speed up the convergence.
+Appendix C Baseline Method
+Backbone.
+The input to our backbone is consecutive thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ resized to 320$\times$256 after data loading, where $T=1$ for the single image setting and $T>1$ for the video setting.
+We initialize the ResNet-18 (he2016deep, ) network with the ImageNet-1K (deng2009imagenet, ) pre-trained weights for the multi-scale features extraction.
+The ResNet-18 (he2016deep, ) backbone returns feature at five levels, with feature sizes as follows given the thermal images as input:
+[…]
+${Features([64,160,128],[64,80,64],[128,40,32],}$
+[…]
+We chose the third level features for future processing considering they carry fine-grained spatial details, which is crucial for accurate mask generation and spatial reasoning in our spatial transformer.
+Mask-guided Spatial Transformer Module.
+For spatial feature refinement, we implement a mask-guided attention mechanism that emphasizes regions of interest within the thermal images, specifically focusing on the hands’ positions.
+This approach utilizes predicted masks from a segmentation head with two convolution layers from the mid-level feature to guide the attention mechanism.
+The dimension of the spatial self-attention module we used is 1024, for which we use an additional one-stride convolution to lift the number of channels of the original mid-level feature.
+Two transformer encoder layers are leveraged and the transformer head number $M$ is set to be 8.
+The deformable attention mechanism introduced in deformable DETR (zhu2020deformable, ) is used and the number of sampling points per attention head $K$ is 4:
+(8) $\small\text{DA}(z_{q},p_{q},x)=\sum_{m=1}^{M}W_{m}\left[\sum_{k=1}^{K}A_{mqk}\cdot W^{\prime}_{m}\left(x(p_{q}+\Delta p_{mqk})\right)\right]$
+where $x$ is the masked low-level feature, $m$ indexes the attention head, $\Delta p_{mqk}$ and $A_{mqk}$ denote the sampling offset and the learnable attention weight for the $k$-th sampling point in the $m$-th attention head respectively.
+Particularly, $A_{mqk}$ is a scalar attention weight lying between 0 and 1, and normalized such that the sum across all $K$ points is 1, $\Delta p_{mqk}$ is a pair of 2D real numbers representing an unconstrained range.
+Noted that 2-D reference point $p_{q}$ is selected from the hand area derived from the predicted mask and $\Delta p_{mqk}$ and $A_{mqk}$ are obtained via linear projection over the query feature $z_{q}$.
+In our self-attention, $z_{q}$ is the added feature for $x$ and the corresponding position encoding.
+Following deformable DETR (zhu2020deformable, ), the first $2M\times K$ channels encode the sampling offsets and the remaining $M\times K$ channels are for the softmax operation to obtain the attention weights.
+The output spatial embedding of the deformable mask-guided spatial attention module has the same size as the input feature, which is [1024, 40, 32].
+Three convolution layers with a stride of 2 are used to downsample the feature map to [128, 10, 8].
+Temporal Transformer Module.
+This module is crucial for understanding temporal patterns over consecutive frames, utilizing a transformer-based approach.
+Given per-frame downsample feature map from the spatial transformer module, we flattened them and apply a linear layer to project this spatial embedding to [1,512].
+Per-frame spatial feature vectors are then fed to the temporal self-attention, which attends to the feature vector of every frame.
+The number of attention layers is $2$ and eight heads are used for each attention.
+The resulting spatio-temporal embedding has the same dimension as the input spatial features, which is [$T$,512].
+Pose Regression. In the pose head, we use the MLP to project the spatio-temporal embedding to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$ for each frame individually.
+$MLP(512\rightarrow 512\rightarrow 512\rightarrow 42\times 3)$
+Loss Functions. Two losses are used for joint learning of hand segmentation and pose estimation during the training process. The L1 hand pose regression loss is defined as:
+(9) $\footnotesize L_{Hand}=\|P^{2D}-P^{2D}_{gt}\|_{1}+\lambda_{1}\|P^{depth}-P^{depth}_{gt}\|_{1}+\lambda_{2}\|P^{3D}-P^{3D}_{gt}\|_{1}$
+where $P^{3D}$, $P^{3D}_{gt}$,$P^{2D}$, $P^{2D}_{gt}$ are the 3D/2D projection of estimated hand joint positions and ground truth hand joint positions.
+$P^{depth}$ and $P^{depth}_{gt}$ are the corresponding depth values for the hand joints.
+$\lambda_{1}$ and $\lambda_{2}$ are the weight parameters set as 100 in our implementation.
+The hand mask binary cross entropy loss is formulated as follows:
+(10) $L_{Mask}$ $=-\frac{1}{WH}\sum_{w=1}^{W}\sum_{h=1}^{H}[w_{pos}\cdot M_{wh}\cdot\log(\hat{M}_{wh})$
+$+w_{neg}\cdot(1-M_{wh})\cdot\log(1-\hat{M}_{wh})]$
+where $M_{wh}$ is the binary ground truth label at pixel location $(w,h)$ and $\hat{M}_{wh}$ is the predicted probability for the estimated mask.
+$w_{pos}$ and $w_{neg}$ are the weights to cope with the foreground-background imbalance.
+We set them to $30:1$ during our training.
+Training.
+We apply the Adam optimizer for our baseline training with an initial learning rate of 1e-4, decayed by a coefficient of 0.95 every epoch.
+All experiments are conducted using two RTX 3090 GPUs, with a batch size of 16 per GPU.
+Models are trained for 25 epochs.

### v5

@@
 Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
-* Equal contribution
-† Corresponding author: xiaoxuan.lu@ucl.ac.uk
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
+3D Hand Pose Estimation, Thermal Vision, Hand Pose Dataset
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 23rd ACM Conference on Embedded Networked Sensor Systems; May 06–09, 2025; Irvine, USAbooktitle: The 23rd ACM Conference on Embedded Networked Sensor Systems (SenSys’25), May 06–09, 2025, Irvine, USAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing → Human computer interaction (HCI)ccs: Computer systems organization → Embedded and cyber-physical systemsccs: Human-centered computing → Ubiquitous and mobile computing
 Figure 1. Data capture setup with the customized head-mounted sensor platform (HMSP) and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
@@
 This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
-Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR images as two ‘spectra’, despite their usual overlap.
+Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap.
 spectral imagery?
-To answer this, we introduce ThermoHands  [^2] Anonymous project page: https://thermohands.github.io/, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging.
+To answer this, we introduce ThermoHands  [^2] Project page: https://thermohands.github.io/., the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging.
 This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig.
@@
   • $\bullet$
-    We introduce the first-of-its-kind benchmark, dubbed ThermoHands, to investigate the potential of thermal imaging for egocentric 3D hand pose estimation.
+    We introduce the first-of-its-kind benchmark, ThermoHands, to investigate the potential of thermal imaging for egocentric 3D hand pose estimation.
   • $\bullet$
@@
   • $\bullet$
-    We will release our dataset, code and models and maintain the benchmark to serve as a new challenge in 3D hand pose estimation.
+    We release our dataset, code and models and maintain the benchmark at https://github.com/LawrenceZ22/ThermoHands.
 2. Related Works
@@
 At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab.
-1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^3] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection..
+1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^3] The study has received the ethical approval from University of Edinburgh, and participant consent forms were signed before the collection..
 As shown in Fig.
@@
 As exhibited in Fig.
-5, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient.
+5, TherFormer features in its two consecutive transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient.
 Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images.
@@
 In practice, our network can be flexibly adapted to the single image-based version (i.e., TherFormer-S) by setting $T=1$.
+Ego-view optimization Multi-view optimization
+Errors $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}+\mathcal{L}_{j3d}$
+mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
+Table 2. Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
+Figure 6. Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
 Mask-guided Spatial Transformer.
@@
 For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately.
-Ego-view optimization Multi-view optimization
-Errors $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}+\mathcal{L}_{j3d}$
-mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
-Table 2. Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
-Figure 6. Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
      Method      Input      MEPE (mm) $\downarrow$      AUC $\uparrow$      MEPE-RA (mm) $\downarrow$      AUC-RA $\uparrow$      fps $\uparrow$
-     (a)      A2J-Transformer (jiang2023a2j, )      Single      51.68      0.474      20.76      0.603      34
-     (b)      HTT (wen2023hierarchical, )      Single      49.09      0.489      20.69      0.599      211
-     (c)      TherFormer-S      Single      44.64      0.539      18.34      0.643      136
-     (d)      (c) w/o spatial transformer      Single      48.79      0.491      20.15      0.609      174
-     (e)      (c) w/o mask guidance      Single      48.83      0.494      18.89      0.625      141
-     (f)      HTT (wen2023hierarchical, )      Sequence      47.07      0.512      17.49      0.659      129
-     (g)      TherFormer-V      Sequence      43.36      0.549      17.36      0.661      52
-Table 3. Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
+     (a)      HaMeR* (pavlakos2024reconstructing, )      Single      -      -      20.88      0.598      118
+     (b)      A2J-Transformer (jiang2023a2j, )      Single      51.68      0.474      20.76      0.603      34
+     (c)      HTT (wen2023hierarchical, )      Single      49.09      0.489      20.69      0.599      211
+     (d)      TherFormer-S      Single      44.64      0.539      18.34      0.643      136
+     (e)      (c) w/o spatial transformer      Single      48.79      0.491      20.15      0.609      174
+     (f)      (c) w/o mask guidance      Single      48.83      0.494      18.89      0.625      141
+     (g)      HTT (wen2023hierarchical, )      Sequence      47.07      0.512      17.49      0.659      129
+     (h)      TherFormer-V      Sequence      43.36      0.549      17.36      0.661      52
+Table 3.
+Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation.
+The fps indicates the number of inference steps that models can run per second.
+HaMeR* estimates 3D hand pose in a root-aligned manner by default; thus, we can only evaluate MEPE-RA and AUC-RA metrics.
 6. Experiment
@@
 As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec.
-4) and analyze the impact for optimization results.
+4) and analyze the impact on optimization results.
 For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames.
@@
 Method and Implementation.
-To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT (wen2023hierarchical, ) and A2J-Transformer (jiang2023a2j, ), and reproduced them on our dataset for experiments.
+To provide sufficient baselines for follow-up works, we selected three state-of-the-art methods in 3D hand pose estimation: HTT (wen2023hierarchical, ), A2J-Transformer (jiang2023a2j, ) and HaMeR (pavlakos2024reconstructing, ), and reproduced them on our dataset for experiments.
 HTT (wen2023hierarchical, ) is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer (jiang2023a2j, ) only works under the single image-based setting.
-For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines.
+HaMeR (pavlakos2024reconstructing, ) is a model-based method that takes a single image as input and reconstructs a MANO-based hand mesh.
+It estimates each hand individually from a cropped image patch centered on the hand, removing the need for explicit localization.
+For comparison, we extract 3D keypoints from its predicted mesh as the estimated pose.
+We use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines.
 We exclude the additional action block used by HTT (wen2023hierarchical, ), focusing solely on hand pose estimation.
@@
 For the RA space, we align the estimated wrist with its groundtruth position before measuring the joint errors.
-In terms of PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
+In terms of PCK, we report the corresponding Area Under the Curve (AUC) over the 0-80mm/50mm error thresholds for the camera/RA space.
 6.3. Thermal Image-based Results
@@
 3.
-TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT (wen2023hierarchical, ) given the same sequential images as input.
+TherFormer-S outforms three competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT (wen2023hierarchical, ) given the same sequential images as input.
 Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec.
@@
 Such results suggest that thermal images can not only serve as supplements in challenging cases but also can be a viable alternative to other spectra in normal conditions for hand pose estimation.
+7. Limitation and Future Work
+As the first exploration in this field, this work has certain limitations that can guide our future work.
+First, our dataset primarily focuses on controlled indoor environment and challenging scenarios for RGB/NIR spectrum.
+Data collection in other challenging scenarios, especially where thermal imaging can be beneficial or limited, is currently scarce.
+In future works, we plan to expand our dataset beyond the current settings to incorporate a wider range of real-world challenges, including grasping hots objects, reflective surfaces to thermal radiation, ambient heating objects, and more outdoor environments.
+Paired data should be collected by including and removing these challenging factors to quantify their impact to thermal imaging-based 3D hand pose estimation.
+Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands.
+Further efforts could be annotating the fine-grained hand action splits and hand-object contact (brahmbhatt2020contactpose, ).
 8. Conclusion
@@
 We believe our foundational endeavour could set the stage for further research in thermal-based 3D hand pose estimation and its wide application.
-Limitation.
-As the first exploration in this field, this work has certain limitations that can guide our future work.
-For example, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands.
-Further efforts could be annotating the fine-grained hand action splits and hand-object contact (brahmbhatt2020contactpose, ).
+Acknowledgments
+This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1) and the grants from the Cisco Research.
 Appendix
@@
     Appendix C introduce more architecture design, parameter and loss details of TherFormer.
-We also provide more visualization in the format of both images and videos for our data, annotation and results on our anonymous project page.
+We also provide more visualization in the format of both images and videos for our data, annotation and results on our project page.
 Appendix A Multi-spectral Hand Pose Dataset
@@
 For the subsequent frames, as we initialize with the results from the last frame, we only optimize for 60 iterations and initialize the learning rate as 0.05 to speed up the convergence.
+Usage Limitations.
+Thanks to the well-established RGB-based tools like (kirillov2023segment, ; mediapipe_hands_2020, ), we can extract accurate 2D hand information, which, combined with depth images, enable MANO-based 3D hand pose annotation.
+However, this approach relies on adequate illumination and minimal occlusion, as hands may appear degraded in RGB or depth images under challenging conditions such as gloves, darkness, and sun glare, and RGB-based tools lack robustness against these factors.
+Moreover, while our annotation pipeline provides accurate 3D hand pose labels, it is impractical for real-time egocentric estimation due to its reliance on multi-view input, high computational cost, and iterative MANO optimization.
+In contrast, 3D hand pose estimation methods are designed for real-time inference from a single image view, more suitable for practical deployment.
 Appendix C Baseline Method